### v1

# Periodic robust robotic rock chop via virtual model control

Abstract
Robotic cutting is a challenging contact-rich manipulation task where the robot must simultaneously negotiate unknown object mechanics, large contact forces, and precise motion requirements. We introduce a new virtual-model control scheme that enables knife rocking motion for robot manipulators, without pre-planned trajectories or precise information of the environment. Motion is generated through interconnection with virtual mechanisms, given by virtual springs, dampers, and masses arranged in a suitable way. Through analysis and experiments, we demonstrate that the controlled robot behavior settles into a periodic motion. Experiments with a Franka manipulator demonstrate robust cuts with five different vegetables, and sub-millimeter slice accuracy from 1 mm to 6 mm at nearly one cut per second. The same controller survives changes in knife shape and cutting board height, and adaptation to a different humanoid manipulator, demonstrating robustness and platform independence.
Index Terms:
Dexterous Manipulation; Compliance and Impedance Control; Robust/Adaptive Control of Robotic Systems; Motion Control of Manipulators.
I Introduction
As a complex manipulation task, robotic cutting requires effective control of the interaction between the cutting tool and its environment, including the geometries of specific objects and their material properties. Contact-rich manipulation is a central theme of robotics research, with significant progress achieved through learning-based approaches [1, 2, 3, 4, 5] and model-based approaches [6, 7, 8, 9]. At the low level of control algorithms, impedance control [10, 11, 12] is often deployed to tame the complex interaction with the environment. However, despite the significant advances in robotic manipulation, robotic cutting remains an open problem. The complex interaction among the knife, the material being cut, and the surrounding environment makes trajectory planning a non-trivial task. Furthermore, determining how to dynamically adjust the manipulator’s impedance in a task-relevant manner is an ongoing area of research.
Consider the ‘trivial’ task of cutting in a kitchen. The knife repeatedly comes into contact with objects of varying material properties, ranging from soft fruits, such as strawberries, to stiffer objects, such as pumpkins and a rigid cutting board. The soft mechanics of food make accurate simulation challenging and this, in turn, complicates the application of learning- and optimization-based methods. In practical settings, learning approaches require a large number of interactions to converge to an effective control policy [13, 14, 15, 16], which not only produces significant food waste but also poses safety risks when experimenting with a knife. Moreover, successful cutting is not just about separating the food; it must also satisfy additional criteria, such as achieving an optimal cut frequency for efficiency and ensuring precise slice thickness to meet culinary presentation standards [17].
Figure 1: Snapshots from robotic cutting.
Accurate trajectory tracking and planning optimization do not seem to yield satisfactory results in the case of cutting. Consider the rocking chop - a technique commonly used in vegetable cutting where the knife is rolled back and forth. Executing such a motion based on controlled predefined trajectories requires precise knowledge of the blade’s geometry to plan a corresponding trajectory [18]. It also requires an exact knowledge of the environment, namely the height and geometry of the cutting board and perhaps of the food. In contrast, it would be highly desirable to have a robust approach that does not depend on precise information, such as the knife geometry. Conventional trajectory generation methods lack this generalization capability and typically rely on accurate vision sensing for re-planning [19, 20, 21].
In this paper, we propose a novel approach that is less reliant on accurate information. Our approach is a virtual model controller (VMC) [22, 23] capable of generating complex motions for contact-rich manipulation tasks. In our approach, the desired robot behavior emerges from the interaction between the robot’s physical mechanics, a virtual mechanism given by an organized collection of mass, springs, and dampers tailored to the task, and the environment. For cutting, these virtual components can realize the necessary motion without requiring explicit trajectory or force tracking. The cutting action remains feasible even if the knife, cutting board or object changes, though the cutting performance may degrade and the resulting motion varies. Moreover, by incorporating sensor feedback, cutting state can be monitored and virtual components can be optimized online to enhance cutting performance such as frequency and force needed for successful separation. Importantly, our controller does not require extensive training data, and tuning is efficient thanks to the intuitive virtual mechanism design.
The central hypothesis of this work is that virtual model control offers a robust solution for interactive tasks, addressing the inherent limitations of trajectory-based methods, specifically their reliance on precise environmental information and the computational burden of adapting to changes. This paper contains the following contributions:
  • a virtual mechanisms controller that realizes rocking motion without explicit trajectory planning;
  • an analytical and experimental demonstration that the controlled robot produces a periodic cutting motion;
  • demonstration of successful cutting of several vegetables, with variable thicknesses and speeds, analyzed through a selection of performance metrics;
  • demonstration of robust cutting, through testing against different environments;
  • real-time adaptation of the virtual mechanism parameters, to improve performance.
The paper is organized as follows. A brief overview of the robotic cutting landscape is summarized in Section II. The methodology of virtual model control and a detailed description of the virtual model controller for cutting are provided in Section III. Section IV provides the details of the implementation and an illustration of the robot motion. An energy-based analysis of the controlled robot behavior can be found in Section V. Experimental benchmarks are discussed in Section VI, with a focus on performance and robustness. Conclusions and further research directions are explored in Section VII, which also provides a detailed comparison with related works.
II Related works
Predicting the mechanical interaction during food cutting is inherently challenging. One of the reasons is that developing an accurate physical model of foodstuffs is complex. The force required to cut through a food item depends on numerous factors, including ripeness, skin presence, orientation, slice thickness, and the knife’s rake angle and sharpness [24, 25, 26, 27]. Simply applying a high cutting force can yield imprecise cuts or even damage delicate produce such as tomatoes. In earlier studies, cutting force was often treated as an external disturbance, with adaptive controllers employing position and velocity feedback to maintain predetermined trajectories [28]. Impedance control was introduced to regulate cutting force directly, augmented by adaptive laws to accommodate time-varying desired positions and environments [29].
Robotic cutting is commonly approached as a motion tracking problem, often informed by visual sensing. A key challenge lies in formulating the cutting trajectory itself. While basic trajectories can be designed, more sophisticated cutting trajectories are harder to derive. Remarkably, [19] obtained a smoothed path from camera data to separate meat muscle. [20] proposed a cut and slice trajectory adjusted through force sensing, to reduce damage to deformable objects. More complex rolling (or rocking) actions have been addressed by leveraging the knife geometry to maintain a constant contact force [18]. [30] explored learning from demonstration, introducing a coupling term via the Udwadia–Kalaba method to suppress lateral blade movements. Reinforcement learning is used in [16] to learn a cutting trajectory and impedance parameters. Because developing an exact physical model of foodstuffs is challenging, learning-based approaches have also garnered attention. A series of works [13, 14, 15] employs data-driven model predictive control (MPC), where a neural network is used to predict future knife positions based on current position and force measurements, thus enabling real-time optimization of the reference force. In [31], variational autoencoders are used to model the cutting dynamics, though their primary goal is to detect material properties and thickness rather than to learn a cutting strategy.
The main difference with our approach is that our controller can generate complex rocking motion without requiring a predefined trajectory. Our controller design and deployment require minimal information about the food and the external environment (knife geometry, cutting board features). Furthermore, its design is not data-intensive and both design and deployment require minimal computing efforts. In turn, these features guarantee the robot achieves successful cutting, robustly handling different food types.
A second challenge involves managing unanticipated interactions with the environment. Inaccurate camera-based detection of the object position, or variations in the table height, can lead to collisions or excessive forces if the robot strictly follows a fixed cutting path. This often results in safety triggers when the commanded position is physically unachievable (e.g., below the cutting board). To deal with the interaction, Yan et al. [32, 18] adopt a hybrid scheme that switches to impedance or direct force control upon detecting contact with the cutting surface. Their method segments the cutting process into distinct press, push, and slice phases, applying different controllers in each stage. In our approach, the controller does not require the identification of different stages of cutting. Rocking is the result of the continuous interaction between robot mechanics, virtual model controller, and environment. Unanticipated interactions may degrade the performance, but only large perturbations will disrupt the whole rocking motion.
A consistent benchmarking framework remains largely absent in the field of robotic cutting. Although prior work often reports end-effector trajectory and force sensor data for individual cutting trials, these metrics are insufficient for real-world applications. In particular, reliability—whether a robot can repeatedly cut and separate multiple slices with accurate thickness—remains understudied. To address this gap, we quantify the performance of our controller against a set of relevant metrics such as cut frequency, cutting force, cutting speed, slice-thickness accuracy and variance.
III Virtual mechanism design for cutting
III-A Virtual model control in a nutshell
A virtual model controller consists of a (virtual) mechanical linkage combined with springs, dampers, and masses that, once coupled with the robot, shapes its behavior. It achieves this by imposing constraints on the motion of the robot and its reaction to the environment. The associated virtual forces are then realized through the robot’s available actuation, typically mapped into joint torques and executed by the motors at each joint [22, 33, 23]. This mapping relies on the robot’s kinematics.
For fully actuated robot manipulator based on revolute joints, using $\boldsymbol{q}$ to denote the joint coordinates of the robot, the force $\boldsymbol{f}_{i}$ generated by a virtual component at a point $\boldsymbol{h}_{i}(\boldsymbol{q})$ of the robot is mapped into corresponding joint torques $\boldsymbol{\tau}$ using the Jacobian matrix $J_{i}(\boldsymbol{q})=\frac{\partial\boldsymbol{h}_{i}(\boldsymbol{q})}{\partial\boldsymbol{q}}$. From the principle of virtual work [34], the torque command to emulate the forces of the virtual model controller is given by
$\boldsymbol{\tau}=J_{1}(\boldsymbol{q})^{T}\boldsymbol{f}_{1}+J_{2}(\boldsymbol{q})^{T}\boldsymbol{f}_{2}+J_{3}(\boldsymbol{q})^{T}\boldsymbol{f}_{3}+\dots$ (1)
Virtual model control is intrinsically stable, because the controlled robot is the interconnection of the passive mechanics of the robot and the passive virtual mechanics [23]. Stability is guaranteed by passivity theory [35, 36, 37, 38, 39, 40, 41]. Virtual model control provides an intuitive, task-oriented method for energy shaping and damping injection in robotics [42, 43].
The selection of spring and damper parameters and their geometric configuration also shapes the impedance of the robot [10, 44]. Although impedance control is typically implemented with linear springs and dampers, a nonlinear spring where force saturates can limit excessively rapid movements and enhance safety. In this work, all springs in the virtual mechanism have a saturating nonlinear characteristic between displacement and force given by
$\boldsymbol{f}(k,\sigma,\boldsymbol{z})=\sigma\tanh(k|\boldsymbol{z}|/\sigma)\frac{\boldsymbol{z}}{|\boldsymbol{z}|},$ (2)
where $k$ is the stiffness parameter, the maximum force is governed by $\sigma\geq 0$, and $\boldsymbol{z}$ represents the spring displacement. Fig. 2 shows the force profile of a spring with stiffness 25 N/m and maximum force 20 N.
Figure 2:  [^20] N saturated nonlinear spring force profile.
III-B Virtual model control for cutting
(a) Rocking motion overview. The knife pivots from tip to heel against the cutting board.
(b) Left: Maintaining contact. A spring and damper link the knife tip $\boldsymbol{p}_{1}$ to $\boldsymbol{p}_{a}$, keeping the blade in contact with the board while being able to slide along the rocking direction ($y$ axis). Middle and right: Cutting and raising. Spring-damper pair connects the knife handle $\boldsymbol{p}_{2}$ to $\boldsymbol{p}_{b}$ which moves vertically for cutting and raising the knife.
(c) Switching references smoothly for cutting (left) or raising (right). Spring-damper pair connects $\boldsymbol{p}_{b}$ to $\boldsymbol{r}_{2,1}$ for cutting. Switching from $\boldsymbol{r}_{2,1}$ to $\boldsymbol{r}_{2,2}$ raises the knife. Instead of directly connecting $\boldsymbol{p}_{2}$ to $\boldsymbol{r}_{2,1}$ or $\boldsymbol{r}_{2,2}$, the mass-spring-damper here produces a smoother transition in $z$ axis. $\boldsymbol{p}_{b}$ slides along a rail. The rail is rigidly connected to $\boldsymbol{r}_{2,1}$ and $\boldsymbol{r}_{2,2}$.
(d) Constraining the slicing plane. Left: Define four fixed points $p_{3},\dots,p_{6}$ relative to the knife. Middle: Define four virtual masses, each can slide along the two axes in the desired slicing plane. Right: Each of the points $p_{3},\dots,p_{6}$ is connected to an individual point mass with a spring and damper.
Figure 3: Virtual mechanism design for rocking motion.
A rocking motion involves continuously shifting the knife’s contact point along the cutting board, from its tip to its heel, as illustrated in Fig. 3(a). To achieve this using a virtual mechanism, we first define a virtual mass that slides along the rocking direction ($y$ axis in Fig. 3(b)) and attach a spring and damper from the knife tip at position $\boldsymbol{p}_{1}$ to the position of the virtual mass $\boldsymbol{p}_{a}$ on the cutting board, ensuring that the blade remains in contact with the surface. The goal of the free-sliding virtual mass is to ensure that the knife can move freely while keeping contact with the board, without interfering with the rocking motion. The virtual mass’s position in the slicing direction $y_{\boldsymbol{p}_{a}}$ follows the mass-spring-damper dynamics:
$m_{1}\ddot{y}_{\boldsymbol{p}_{a}}+\boldsymbol{f}(k_{1},\sigma_{1},y_{\boldsymbol{p}_{a}}-y_{\boldsymbol{p}_{1}})+c_{1}(\dot{y}_{\boldsymbol{p}_{a}}-\dot{y}_{\boldsymbol{p}_{1}})=0$ (3)
where $y_{\boldsymbol{p}_{1}}$ is the position of $\boldsymbol{p}_{1}$ in the slicing direction. $c_{1}$ is the damping coefficient.
Another spring and damper connect the knife handle at $\boldsymbol{p}_{2}$ to the control point $\boldsymbol{p}_{b}$ for the cutting motion. Switching $\boldsymbol{p}_{b}$ from reference $\boldsymbol{r}_{2,1}$ to reference $\boldsymbol{r}_{2,2}$ raises the knife. Switching from $\boldsymbol{r}_{2,2}$ to $\boldsymbol{r}_{2,1}$ cuts down (Fig. 3(b)). Hence, the system alternates between a cutting state and a knife-raising state, producing the rocking action. The motion is in cutting phase when $\boldsymbol{p}_{b}$ is connected to $\boldsymbol{r}_{2,1}$ and raising phase when connected to $\boldsymbol{r}_{2,2}$.
Positions $\boldsymbol{p}_{1}$ and $\boldsymbol{p}_{2}$ are defined relative to the robot’s end-effector, assuming that the knife grip is rigid enough, so that the relationship between end-effector and knife does not change. $\boldsymbol{p}_{a}$, $\boldsymbol{r}_{2,1}$ and $\boldsymbol{r}_{2,2}$ lie in the world frame and depend on the position of the food. The difference in $y$ coordinate of $\boldsymbol{r}_{2,1}$ and $\boldsymbol{r}_{2,2}$ regulates the slicing distance. The cutting force is regulated through the stiffness parameters $k_{1}$ and $k_{2}$. The position of $\boldsymbol{p}_{a}$ and the selection of the reference $\boldsymbol{r}_{2,1}$ also affect the cutting force.
The switch from cutting to raising is triggered by the knife’s orientation. When $z$ coordinates of $\boldsymbol{p}_{1}$ and $\boldsymbol{p}_{2}$ are close (within tolerance $\delta_{1}$), the knife spine (the line that passes $\boldsymbol{p}_{1}$ and $\boldsymbol{p}_{2}$) aligns with the cutting board surface and the cut is done. The switch from raising to cutting is triggered when $\boldsymbol{p}_{2}$ is close to $\boldsymbol{r}_{2,2}$ (within tolerance $\delta_{2}$ measured along the z coordinates), to ensure enough height to cut the food.
All together, these components produce virtual forces
$\boldsymbol{f}_{1}$ $=\boldsymbol{f}(k_{1},\sigma_{1},\boldsymbol{p}_{a}-\boldsymbol{p}_{1})+c_{1}(\dot{\boldsymbol{p}}_{a}-\dot{\boldsymbol{p}}_{1})$ (4a)
$\boldsymbol{f}_{2}$ $=\boldsymbol{f}(k_{2},\sigma_{2},\boldsymbol{p}_{b}-\boldsymbol{p}_{2})+c_{2}(\dot{\boldsymbol{p}}_{b}-\dot{\boldsymbol{p}}_{2}).$ (4b)
where $\boldsymbol{p}_{b}$ is switched between $\boldsymbol{r}_{2,1}$ and $\boldsymbol{r}_{2,2}$ depending on whether it is in the cutting or raising phase. The way $\boldsymbol{p}_{b}$ switches is also modulated by the dynamics in Fig. 3(c). We smooth the transition by placing a mass that slides along the line between $\boldsymbol{r}_{2,1}$ and $\boldsymbol{r}_{2,2}$. The mass is connected to $\boldsymbol{p}_{2}$ and $\boldsymbol{r}_{2}$ with springs and dampers.
Finally, to ensure the knife moves primarily within the plane that defines the slice, we introduce four additional points $\boldsymbol{p}_{3},\dots,\boldsymbol{p}_{6}$ on the knife (Fig. 3(d)) and connect each to a virtual mass free to slide along the desired slicing plane, whose position is denoted by the reference $\boldsymbol{r}_{S}$ Each pair is linked by its own spring and damper. This virtual mechanism enforces a plane-constrained motion, yielding a cleaner cut. To produce multiple slices in sequence, once a slice is completed and the knife has returned to its raised position, we shift the slicing plane $\boldsymbol{r}_{S}$, $\boldsymbol{p}_{a}$, $\boldsymbol{r}_{2,1}$, and $\boldsymbol{r}_{2,2}$ to the next desired location.
IV Experimental platform and cutting motion
IV-A Robot platform and control architecture
The virtual model controller outlined in the previous section is implemented through the VMRobotControl.jl package [45], a platform-independent framework designed for virtual model control in robotic systems. This package enables the construction and simulation of both the robot and virtual components. Given the robot’s URDF representation, the package offers several primitives to define the virtual model controller, to compute its dynamics, and to generate the torque commands for each joint.
Our experimental platform consists of a 7-DoF Franka Research 3 arm, with a maximum payload of 3 kg. To collect force data, a ROBOTIQ FT 300-S force-torque sensor is mounted at the robot wrist, providing force and torque measurements at 100 Hz with a noise level of approximately 0.1 N. A knife is rigidly connected to the sensor using a 3D-printed attachment as shown in Fig. 4. Thanks to the platform independence of the virtual model design, the cutting mechanism was also implemented on a humanoid robot. Sciurus17, a lightweight and low-cost platform developed by RT Corp. Scirus17 features 17 degrees of freedom (DoF) distributed across its upper body (with 7 DoF per arm, 2 DoF for the head and 1 DoF for the torso). The robot utilizes Dynamixel motors with current control, and the conversion from torque to current is approximated by means of a linear relation based on the performance characteristics of the motor.
Figure 4: Experiment setup: robot arm with force sensor and knife rigidly attached.
The rocking motion was initially validated via simulations. However, because of the computational cost and potential inaccuracies associated with simulating contact forces for cutting, further tuning was performed during real-world experiments. The controller runs on a standard computer. The overall control architecture is depicted in Fig. 5. Virtual forces $\boldsymbol{f}$ generated by the virtual mechanisms are computed and translated into joint torques $\boldsymbol{\tau}$ using (1). The computation of the current state of the virtual mechanisms, such as the elongation of virtual springs and the velocities used by the virtual dampers, is based solely on the robot’s joint position $\boldsymbol{q}$ and velocity $\dot{\boldsymbol{q}}$. All force computations and their conversion into joint torques rely on forward kinematics and the associated Jacobians. This represents a significant departure from traditional motion planning, where execution is typically based on inverse kinematics.
Figure 5: Control architecture implemented with VMRobotControl.jl package.
IV-B Emergence of cutting motion in experiments
(a) Snapshots from experiment: rocking a courgette with a knife.
(b) Top: Solid: $\boldsymbol{p}_{1}$, Dashed: $\boldsymbol{p}_{a}$. Bottom: Solid: $\boldsymbol{p}_{2}$, Dashed: $\boldsymbol{p}_{b}$.
(c) End effector position and force reading.
(d) End effector orientation in Euler angles and torque reading.
Figure 6: Rocking motion as a result of the interaction between the virtual mechanism and environment.
The desired rocking motion is not predetermined but emerges from the coordinated action of the several parts of the virtual mechanism in Fig. 3. The virtual mechanisms constrain the motion of the blade by enforcing (compliant) contact points and ‘soft’ constraints. Switching of the reference at specific points induces a cyclical motion of the knife, which is also modulated by the interaction among robot dynamics, compliant virtual mechanism constraints, and contacts with the environment. A first illustration is provided in Fig. 6(d), whose control parameters are specified in Table I, A-1. Using courgette as the cutting target, the snapshots in Fig. 6(a) illustrate the continuous contact between the knife and the cutting board while the knife rolls back and forth. The successful separation of the courgette is also seen in the snapshots. Since the motion is not pre-programmed, variations in the knife geometry or cutting board features will yield different motions.
Fig. 6(b) shows the position of two springs responsible for the rocking motion. The sliding of the virtual mass ensures that $\boldsymbol{p}_{a}$ stays close to $\boldsymbol{p}_{1}$ and no additional force is applied in the rocking direction. The displacement between $\boldsymbol{p}_{1}$ and $\boldsymbol{p}_{a}$ along the $z$ axis pulls the knife towards the cutting board and ensures contact. The knife moves to the next slice position when raised. Fig. 6(c) and 6(d) show the end-effector’s position and force, orientation and torque. The resulting end-effector’s motion is a fairly complicated trajectory, difficult to design manually. One would need to know the geometry of the knife to decide where the end-effector should be while rolling the knife back and forth.
Force and torque are measured at the robot’s wrist and transformed to the world coordinate. $F_{y}$ and $F_{z}$ correspond to the forces applied to the robot wrist to slice and cut. $F_{x}$, $\tau_{y}$ and $\tau_{z}$ act to stabilize the knife on the slicing plane. The pitch and yaw remain small, which shows that the stabilization of the slicing plane is effective. The roll and torque $\tau_{x}$ account for the rocking motion. During the lifting of the knife, a higher amount of torque $\tau_{x}$ is applied when the contact point is shifted further to the tip.
V Energy-based analysis of the periodic motion
Energy considerations provide insight into the emerging periodic motion. The mechanical energy of the robot, $E_{R}$, satisfies the following standard passivity inequality [46, 34, 43]
$\dot{E}_{\mathrm{R}}\leq\dot{\boldsymbol{q}}^{T}\boldsymbol{\tau}+\sum_{i}\dot{\boldsymbol{s}}_{i}^{T}\boldsymbol{d}_{i}$ (5)
where $\boldsymbol{q}$ is the vector of joint coordinates of the robot manipulator, $\boldsymbol{\tau}$ are the related motor torques, and $(\boldsymbol{s}_{i},\boldsymbol{d}_{i})$ are additional co-located positions/forces variables, accounting for generic external actions to the robot. The inequality takes into account unavoidable losses due to internal mechanical dissipation.
The controller is also a virtual mechanism, whose mechanical energy satisfies
$\dot{E}_{\mathrm{VMC}}=-\sum_{i=1}^{6}\dot{\boldsymbol{p}}_{i}^{T}\boldsymbol{f}_{i}-W_{\mathrm{dissipation}}+W_{\mathrm{elastic}}.$ (6)
$W_{\mathrm{dissipation}}\geq 0$ accounts for the power dissipated in the dampers of the virtual mechanism. $W_{\mathrm{elastic}}$ captures the power flow caused by the switching of the references $\boldsymbol{r}_{2,1}$, $\boldsymbol{r}_{2,2}$, and $\boldsymbol{r}_{S}$.
The negative sign of the term $-\sum_{i=1}^{6}\dot{\boldsymbol{p}}_{i}\boldsymbol{f}_{i}$ in (6) is due to the mechanical interconnection between robot and virtual model controller. Through (1), each force $\boldsymbol{f}_{i}$ acting on the robot exerts an equal and opposite force on the virtual mechanism. The total energy of the controlled robot, $E=E_{\mathrm{R}}+E_{\mathrm{VMC}}$, thus satisfies
$\dot{E}\leq-W_{\mathrm{dissipation}}+W_{\mathrm{elastic}}+\sum_{i}\dot{\boldsymbol{s}}_{i}^{T}\boldsymbol{d}_{i}$ (7)
(7) shows that the energy of the controlled must continuously decrease if no external forces act on the robot (i.e. $\boldsymbol{d}_{i}=0$) and if we do not switch references $\boldsymbol{r}_{2,1}$, $\boldsymbol{r}_{2,2}$, and $\boldsymbol{r}_{S}$ of the virtual mechanism. The robot motion eventually settles in a resting position. For the gravity-compensated robot, this corresponds to a minimum of the elastic potential energy. This is a classical result of passivity-based control [42].
In our case, each time a reference is switched at a specific robot configuration, a bounded amount of energy is either injected into or removed from the system. This action continuously drives the controlled robot into motion. Switching between the cutting and raising phases is governed by different triggers: the transition from raising the knife to cutting is position-based, while the switch from cutting to raising is primarily triggered by the knife’s orientation. However, because the cutting board restricts the motion of the knife, we can also treat the latter as a position-based trigger. When a switch occurs, the spring that was previously connected to the reference $\boldsymbol{r}_{2,1}$ is detached and reattached to $\boldsymbol{r}_{2,2}$, and vice versa. Because of these reference switches, the virtual springs generate new forces, leading to non-zero velocities and thus new dissipation ($W_{\mathrm{dissipation}}>0$) in the virtual dampers. The injected elastic energy is always bounded, due to the position-based switching mechanism.
The periodic cutting motion of the controlled robot results from the balance between the bounded energy injected through reference switching and the energy dissipated by virtual dampers, the internal dissipation of the robot, and the environment. That is, neglecting internal dissipation and environment contributions, for any periodic motions of period $T$ we have, for all $t$,
[…]
The overall system does not accumulate energy indefinitely. The injected energy is either dissipated by the robot’s inherent damping and the virtual dampers, or transferred into the environment (e.g., performing cutting work on the vegetable or contacting the board). Hence, over multiple cycles, the system settles to a bounded regime of energy. Any surplus energy in the system regardless of its source ultimately appears as increased speed at switching or larger cutting forces. Nonetheless, it is dissipated through robot’s physical and virtual damping, or absorbed by subsequent contact with the board or vegetable, preventing unbounded energy growth.
Further insights can be gathered by looking at a toy model, given by a single mass and one degree of freedom. For a single controlled mass we have
$m\ddot{x}+c\dot{x}+k(x-r)=0,$ (9)
where $m>0$ is the mass, $c>0$ is the damping coefficient, and $k>0$ is the spring stiffness. $r$ is the reference, switching between two values, $r\in\{r_{1},r_{2}\}$. $x$ is the mass position. The mechanical energy reads
$E=\frac{1}{2}m\dot{x}^{2}+\frac{1}{2}k(x-r)^{2}$ (10)
from which
$\dot{E}=-c\dot{x}^{2}+W_{r}.$ (11)
$W_{r}$ accounts for the flow of energy due to switching reference. For instance, $W_{r}$ is zero unless a switch occurs. A switch at time $t$ causes an instantaneous (and large!) flow of energy, $W_{r}\neq 0$. The amount of injected energy can be computed by looking at the left and right limits of the switching time $t$, denoted by $t^{-}$ and $t^{+}$. For instance,
${E(t^{+})-E(t^{-})}=\frac{1}{2}k(x(t^{+})-r(t^{+}))^{2}-\frac{1}{2}k(x(t^{-})-r(t^{-})^{2}.$ (12)
If $r\in\{r_{1},r_{2}\}$ and switch times depend on $x$, this injection is always bounded. A periodic motion arises at the balance between injection of energy (switches) and dissipation (during motion). Consider the case that a switch occurs at positions $x_{1}$ and $x_{2}$, as depicted in Fig. 7. This case mirrors the cutting/raising transitions in rocking motion. Formally,
$r$ $=r_{1}\quad\text{if}\quad\bigl{(}\dot{x}\geq 0\,\wedge\,x\leq x_{1}\bigr{)}\;\text{or}\;\bigl{(}\dot{x}\leq 0\,\wedge\,x\leq x_{2}\bigr{)},$
$r$ $=r_{2}\quad\text{if}\quad\bigl{(}\dot{x}\geq 0\,\wedge\,x\geq x_{1}\bigr{)}\;\text{or}\;\bigl{(}\dot{x}\leq 0\,\wedge\,x\geq x_{2}\bigr{)}.$
Figure 7: Switching the reference between $r_{1}$ and $r_{2}$ at positions $x_{1}$ and $x_{2}$. This schematic mirrors the cutting/raising transitions in rocking motion.
By plotting trajectories in the $(x,\dot{x})$ phase plane (Fig. 8), one observes that solutions converge to a closed curve. Intuitively, if the mass moves too fast, the damper dissipates more energy, returning the system toward this closed curve. Conversely, if motion is too slow, less energy is dissipated, and the system gains more momentum until it approaches the same closed curve.
A final analysis can be developed experimentally. Consider a generic cutting cycle identified by the robot switching reference from ‘raising’ to ‘cutting’. The total work done by the robot per cutting cycle can be approximated by the energy
$E_{\mathcal{C}_{j}}\,=\,\sum_{t_{i}\in\mathcal{C}_{j}\setminus\{t_{N}\}}\boldsymbol{F}\bigl{(}t_{i}\bigr{)}\bigl{(}\boldsymbol{x}(t_{i+1})-\boldsymbol{x}(t_{i})\bigr{)},$ (13)
where $\boldsymbol{F}$ is the force read from the force sensor placed at the end-effector while cutting. $\boldsymbol{x}$ is its position. The set $\mathcal{C}_{j}=\{t_{1},\dots,t_{N}\}$ contains the time instants of cycle $j$ at which sample measures are taken. At each cycle, $E_{\mathcal{C}_{j}}$ provides an estimate of the amount of work required to cut, considering all dissipations and interactions (internal, external, food, cutting board).
We start the experiment with the knife initially performing a pure rocking motion with no food. From the 7th to the 16th cycle, a carrot is introduced and cut by the knife, after which the carrot is removed. As illustrated in Fig. 9, the work done for cutting the carrot differs from the work done during baseline rocking motion. However, once the carrot is removed, the system returns to the energy level observed during pure rocking. This behavior shows how the virtual model controller adjusts the robot motion to compensate for external perturbations, and consistently returns to prior energy balance when perturbations are removed.
Figure 8: Phase plane of a mass spring damper system with switching references $r_{1}=1$ and $r_{2}=-1$ and positions $x_{2}=-0.8$ and $x_{1}=0.8$. Trajectories with different initial conditions converge to the same limit cycle.
Figure 9: Work done for each cutting cycle. Red: cutting a carrot, blue: no food
VI Experimental result and discussion
VI-A Forces, velocities, and frequency of successful cut
(a) Black: Force reading in the z axis. Red: Switch from raising to cutting. Blue: Switch from cutting to raising. Dash: Same control parameters for all food. Continuous: Tuned control parameters for each food so that the knife cuts through without applying excessive force.
Figure 10: Rocking mechanism applied to five different food. Left two columns: blue - untuned, red - tuned for minimal force.
We evaluated the versatility of the controller by cutting a variety of vegetables (spring onion, cucumber, carrot, courgette, and potato) under two conditions: (i) using a single set of control parameters for all food (uniform parameters), and (ii) with parameters tuned for each food type (food-tuned parameters), with the goal of minimizing the applied force. In fact, reducing the cutting force helps to achieve precise cuts and prolongs the lifespan of both the knife and the cutting board. Larger items, such as cucumber, courgette, and potato, were cut in half to make them suitable for rocking and ensure stable placement on the cutting board. All vegetables were securely held to prevent motion during cutting. Results are summarized in Fig. 10.
To quantitatively evaluate the cutting process, we have used the following metrics. Let $\mathcal{C}_{\mathrm{cut},i}=\{t_{1},\dots,t_{N_{i}}\}$ denote the times at which a sample measure is taken during the cutting phase $i$. These belong to the time interval in which $\boldsymbol{p}_{b}$ is connected to the reference $\boldsymbol{r}_{2,1}$, as shown in Fig. 3(c) (left). The average and peak forces during the $i$th cycle are computed from the force-torque sensor reading, mounted at the robot wrist
$F_{\text{avg},i}\,=\!\!\sum_{t\,\in\,\mathcal{C}_{\mathrm{cut},i}}\!\!\frac{|\boldsymbol{F}(t)|}{N_{i}}\qquad\quad F_{\text{peak},i}\,=\max_{t\,\in\,\mathcal{C}_{\mathrm{cut},i}}|\boldsymbol{F}(t)|.$ (14)
Similarly, the average and maximum velocities of the robot wrist during cutting are estimated as follows.
$v_{\text{avg},i}\,=\!\!\sum_{t\,\in\,\mathcal{C}_{\mathrm{cut},i}}\!\!\frac{|\boldsymbol{v}(t)|}{N_{i}}\qquad\quad v_{\text{peak},i}\,=\max_{t\,\in\,\mathcal{C}_{\mathrm{cut},i}}|\boldsymbol{v}(t)|,$ (15)
where $\boldsymbol{v}$ is the velocity of the wrist, computed using the robot kinematics. Finally, cut frequency is defined as
$f_{\text{cut}}\,=\,\frac{N_{\text{switch}}-1}{T},$ (16)
where $N_{\text{switch}}$ is the total number of transitions from the cut to the raise state that occur during the time interval $T$ of a whole experiment. This frequency indicates how rapidly consecutive slices are produced. Work for one slice is computed with (13).
When using a single set of control parameters for all foods, the force profiles in the $z$-axis (Fig. 10(a)) were similar across all food types despite significant differences in stiffness. This is because the control parameters were chosen to apply a sufficiently high force so that the intrinsic properties of the food had minimal impact on the resulting behavior. Peak force is typically observed near the transition from cut to raise and varies with the energy absorption characteristics of each vegetable. The level of ‘residual’ energy after cut typically results in a higher contact force before the knife is retracted. With parameters tuned specifically for each food type, the force required for cutting spring onion and cucumber is noticeably reduced, while high force levels are needed for other vegetables. Fig. 1 shows the sliced vegetables at the end of experiments. By visual inspection, both (i) uniform and (ii) food-specific parameter selection produce no noticeable difference in slice quality or accuracy.
Figure 11: Cutting different slices. Top row: blue - 0.94 cut/sec, red - 0.51 cut/sec. Bottom row: fast and slow frequencies.
VI-B Cut accuracy
To evaluate the precision of our cutting mechanism, we cut courgette samples using various target thicknesses and cutting speeds. Two average cut frequencies were implemented for each thickness setting, approximately 0.51 and 0.94 cuts per second. Fig. 11 shows representative slices obtained under these conditions.
Cutting accuracy is assessed by measuring each cut in four distinct locations. Measurements are taken with care, avoiding deformation of the slice as much as possible. We use three metrics to measure (i) if the desired thickness is achieved, (ii) if the accuracy is consistently achieved with continuous cutting, and (iii) if each slice is evenly cut.
For (i), we compare the average of all measurements $\Delta$ with the desired one. We compute
$\Delta=\frac{1}{N}\sum_{i=1}^{N}\Delta_{i}\qquad\quad\Delta_{i}=\frac{1}{4}\sum_{j=1}^{4}\Delta_{i,j}$ (17)
where $N$ is the total number of slices, $\Delta_{i}$ is the average thickness of slice $i$ based on four thickness measurements $\Delta_{i,j}$. For (ii), we take the variance of the average thickness slices. Namely,
$\sigma^{2}_{\Delta}=\frac{1}{N}\sum_{i=1}^{N}\left(\Delta_{i}-\Delta\right)^{2}.$ (18)
Higher variance indicates lower consistency. Finally, for (iii), we take the variance of the four thickness measurements within each slice and compute its average across all slices:
$\sigma^{2}_{\Delta}=\frac{1}{N}\sum_{i=1}^{N}\sigma^{2}_{\Delta_{i}}\qquad\sigma^{2}_{\Delta_{i}}=\frac{1}{4}\sum_{j=1}^{4}\left(\Delta_{i,j}-\Delta_{i}\right)^{2}.$ (19)
In this case, a higher variance indicates lower evenness.
Fig. 11 shows that average thickness is close to the desired target. In particular, the slices are consistently slightly thinner than expected. Several factors may contribute to this discrepancy: the knife’s movement matches the desired thickness without compensating for its own width, alongside the plastic deformation of the vegetable during cutting and moisture loss in the slices. Despite these effects, achieving nearly one slice per second with such precision is comparable to or even exceeds human performance.
Faster cutting generally degrades performance in terms of average thickness, consistency, and evenness, as illustrated in Fig. 11, except in the case of 1-mm slices. The decreased performance at higher speeds is mainly due to the limited ability of the knife to accurately reach the target slicing position in a short time frame, resulting in fluctuations and positional adjustments during the cut. This effect is minimized when cutting very thin slices. Additionally, 1-mm slices experience greater plastic deformation at slower cut speeds, potentially hindering the initiation and propagation of a clean cut.
VI-C Cut robustness
The VMC approach demonstrates inherent robustness in handling environmental variations. As shown in Section VI-A, the same set of control parameters enabled successful cut of a variety of food items, from delicate spring onions to rigid potatoes. In addition to these results, we tested the controller using three different knives, as shown in Fig. 12(a). Knife 1 features an optimal blade curvature for rocking and has been used in all experiments so far. Knife 2 has a different blade geometry, which is still suitable for rocking. In contrast, knife 3 is a carving knife and should not be used for rocking. Results are shown in Fig. 12. To assess robustness, we consider a uniform set of parameters for all knives.
Although all knives exhibited rocking motion under identical control parameters, the differences in blade geometry affected energy dissipation and resulted in different rocking features. Knife 2 had a smaller blade than knife 1, resulting in a reduced energy dissipation through contact with the cutting board. This led to a higher residual energy after cutting, which turned into speed, that is, faster cuts. Similar considerations apply to knife 3. The progressive increase in force peaks for knives 2 and 3, respectively, can be explained in a similar way: knives 2 and 3 are moving faster than knife 1, leading to stronger forces at contact with the cutting board.
We also demonstrate the ability of the controller to handle different heights of the cutting board. Fig. 13(a) illustrates the experimental setup, where the nominal cutting board height was modified by adding MDF layers underneath. Fig. 13 shows that the rocking motion remained stable across all heights, without requiring any adjustments to the control parameters. The gradual decrease in peak force with increasing height can be explained by the knife having less distance to accelerate before switching.
Together, Figs. 12 and 13 provide a neat illustration of the difference between our virtual model controller and trajectory-based methods. The latter would require a precise knowledge of knife geometry and new path planning computations to take into account variations in cutting board heights. In contrast, because the rocking motion established by the virtual model controller is not programmed, it can easily comply with various settings.
(a) Three different knives.
(b) Knives 1, 2, and 3 rocking motions.
(c) Forces, velocities, and cut frequencies of the three knives.
Figure 12: Rocking motion with three different knives while using the same virtual mechanism and control parameters.
(a) Experiment setup with different cutting board heights.
(b) Forces and velocities for different heights of the board.
Figure 13: Rocking with different cutting boards.
VI-D Platform independence
The proposed virtual mechanism is platform independent. We demonstrate this by implementing the same controller on a humanoid robot from RT Corp, Sciurus17. Unlike the Franka platform, Sciurus17 has a lower payload capacity and slower communication speed. This can be managed by re-tuning the parameters of the controller to ensure smaller forces and slower motions. Fig. 14 shows that the inherent limitations of the robotic platform led to reduced performance. However, successful rocking motion is achieved.
(a) Snapshots of Sciurus17 rocking with a knife.
(b) Top: solid - $\boldsymbol{p}_{1}$, dashed - $\boldsymbol{p}_{a}$. Bottom: solid - $\boldsymbol{p}_{2}$, dashed - $\boldsymbol{p}_{b}$.
Figure 14: Rocking motion Sciurus17. The virtual model controller is unchanged but its parameters have been adjusted.
VII Conclusion and future works
We have proposed a new virtual model controller for rock cutting. The rocking motion has been discussed theoretically, through energy considerations, and experimentally. Performance has been illustrated by cutting several vegetables, at several thicknesses and cut frequencies. Robustness has been illustrated by using three different knives and several cutting board heights. Notably, the controller does not require any trajectory planning and relies on minimal information, both at design and deployment stage, and delivers consistent and reliable cutting performance. Experiment videos are available online: https://youtu.be/67v-U4gp9_k.
Although these results clarify the flexibility and reliability of our control strategy, there are still significant directions for improvement. At level of integration, the most pressing one is handling the full cutting pipeline. This entails autonomous pick and place of the food, its orientation on the cutting board, and the capability of holding food in place during cutting. It seems also crucial for the robot to recognize different foods and knives, and to measure in real-time the result of its action, during and after cutting. More work is also needed at the level of the food-knife interaction. A key future direction is the systematic exploration of adaptation/learning of the control parameters, to improve performance. The idea is to exploit additional sensing, such as vision and tactile information, to optimize control parameters in real time.
Thanks to the intuitive design of our virtual mechanism, automatic control parameter adjustment can be performed at different levels of complexity. Here, for completeness, we discuss an initial study in which the spring stiffness $k_{2}$ is adjusted to achieve successful food separation while moderating the force exerted on food and board. The overall strategy is extremely simple. We increase $k_{2}$ if the knife does not achieve separation of food (by position sensing through robot and knife geometry). We reduce $k_{2}$ if we detect excess overshoot on the knife handle (i.e. if the knife handle moves past the critical reference). This leads to an error signal, $e$, capturing the mismatch between desired and achieved vertical position of the knife. We update $k_{2}$ every six cutting cycles, using the adaptation law
$k_{2}^{(n+1)}=k_{2}^{(n)}+\alpha\,e,$
where $\alpha$ is the adaptation gain.
Figure 15 provides a brief illustration. All control parameters are set as in the A-2 Fast cut in Table I, except for $k_{2}$, which is dynamically adapted. $k_{2}$ starts from a low value, which is sufficient initially to cut through a carrot. From time $t=20$s, the carrot presents a harder consistency which requires adaptation to cut through as shown in Figure 15(b). $k_{2}$ increases and, after $t=35$s, adaptation is completed and cutting is reestablished. This brief illustration shows the potential of continuous learning combined with virtual model control. Here we have focused on a simple single parameter for simplicity, but we plan to develop a systematic approach in the future, leveraging vision and tactile sensing.
(a) Carrot cut result.
(b) Stiffness $k_{2}$.
Figure 15: Adaptation of spring stiffness for cutting a carrot.
TABLE I: Control parameters for each experiment. The following parameters remain the same across all experiments: $m_{a}=0.1$, $m_{\text{ori}}=0.1$.
(a) Performance
Exp $k_{1}$ $\sigma_{1}$ $c_{1}$ $k_{2}$ $\sigma_{2}$ $c_{2}$ $k_{\text{ori}}$ $\sigma_{\text{ori}}$ $c_{\text{ori}}$ $\boldsymbol{r}_{2,1}\,(y,z)$ $\boldsymbol{r}_{2,2}\,(y,z)$ $\delta_{1}$ $\delta_{2}$ $m_{b}$ $k_{b}$ $c_{b}$
A-1 Uniform parameters 25 10 1 150 25 2 1200 50 10 0.43, -0.1 0.58, 0.4 0.02 0.15 1 30 17
A-1 Spring onion 10 10 1 65 25 2 1200 50 10 0.43, -0.1 0.58, 0.4 0.02 0.2 0.01 30 1.7
A-1 Cucumber 10 10 1 75 25 2 1200 50 10 0.43, -0.1 0.83, 0.5 0.02 0.25 0.01 30 1.7
A-1 Courgette 10 10 1 105 25 2 1200 50 10 0.43, -0.1 0.83, 0.5 0.02 0.25 0.01 30 1.7
A-1 Carrot 20 10 1 105 25 2 1200 50 10 0.43, -0.1 0.83, 0.5 0.02 0.25 0.01 30 1.7
A-1 Potato 20 10 1 105 25 2 1200 50 10 0.43, -0.1 0.83, 0.5 0.02 0.25 0.01 30 1.7
(b) Performance
Exp $k_{1}$ $\sigma_{1}$ $c_{1}$ $k_{2}$ $\sigma_{2}$ $c_{2}$ $k_{\text{ori}}$ $\sigma_{\text{ori}}$ $c_{\text{ori}}$ $\boldsymbol{r}_{2,1}\,(y,z)$ $\boldsymbol{r}_{2,2}\,(y,z)$ $\delta_{1}$ $\delta_{2}$ $m_{b}$ $k_{b}$ $c_{b}$
A-2 Fast cut (0.94 cut/s) 45 10 1 200 30 2 800 50 10 0.43, -0.1 0.73, 0.35 0.035 0.145 0.01 250 4
A-2 Slow cut (0.51 cut/s) 25 10 1 120 25 2 800 50 10 0.43, -0.1 0.58, 0.4 0.02 0.15 1 30 17
(c) Robustness
Exp $k_{1}$ $\sigma_{1}$ $c_{1}$ $k_{2}$ $\sigma_{2}$ $c_{2}$ $k_{\text{ori}}$ $\sigma_{\text{ori}}$ $c_{\text{ori}}$ $\boldsymbol{r}_{2,1}\,(y,z)$ $\boldsymbol{r}_{2,2}\,(y,z)$ $\delta_{1}$ $\delta_{2}$ $m_{b}$ $k_{b}$ $c_{b}$
B Robustness 30 10 1 130 25 2 800 50 10 0.43, -0.1 0.63, 0.4 0.01 0.2 0.005 30 3
(d) Platform independence, Sciurus
Exp $k_{1}$ $\sigma_{1}$ $c_{1}$ $k_{2}$ $\sigma_{2}$ $c_{2}$ $k_{\text{ori}}$ $\sigma_{\text{ori}}$ $c_{\text{ori}}$ $\boldsymbol{r}_{2,1}\,(y,z)$ $\boldsymbol{r}_{2,2}\,(y,z)$ $\delta_{1}$ $\delta_{2}$ $m_{b}$ $k_{b}$ $c_{b}$
C Sciurus 10 2 0.3 20 10 0.1 30 10 1 0.4, -0.1 0.6, 0.6 0.047 0.38 0.8 110 50