### v2

# Periodic robust robotic rock chop via virtual model control

Abstract
Robotic cutting is a challenging contact-rich manipulation task where the robot must simultaneously negotiate unknown object mechanics, large contact forces, and precise motion requirements. We introduce a new active virtual-model control scheme that enables knife rocking motion for robot manipulators, without pre-planned trajectories or precise information of the environment. Motion is generated and controlled through switching virtual coupling with virtual mechanisms, given by virtual springs, dampers, and masses arranged in a suitable way. Through analysis and experiments, we demonstrate that the controlled robot behavior settles into a periodic motion. Experiments with a Franka manipulator demonstrate robust cuts with five different vegetables, and sub-millimeter slice accuracy from 1 mm to 6 mm at nearly one cut per second. The same controller survives changes in knife shape and cutting board height, and adaptation to a different humanoid manipulator, demonstrating robustness and platform independence.
keywords:
Dexterous Manipulation; Compliance and Impedance Control; Robust/Adaptive Control; Motion Control.
1 Introduction
Robotic cutting captures the complexity of contact-rich tasks. Cutting requires dexterous, precise motion that adapt to different surfaces, knife geometries, and food mechanics. Contact-rich manipulation is a central theme of robotics research, with significant progress achieved through learning-based approaches Popov et al. (2017); Rajeswaran et al. (2018); Gupta et al. (2021); Radosavovic et al. (2021); Chen et al. (2022) and model-based approaches Posa et al. (2014); Marcucci et al. (2017); Hogan and Rodriguez (2020); Aydinoglu and Posa (2022). At the level of control algorithms, contact-rich interactions require impedance control Hogan (1985a, b, c). Advanced control over the force-displacement relationship at the end-effector is essential to regulate complex interactions with the environment. However, despite significant progress in robotic manipulation, autonomous cutting remains an open challenge due to the unpredictable nature of material fracture and deformation. In addition, the specific geometries of the knife and cutting surface are often only partially known. This uncertainty in the mechanical coupling between the tool and the workspace prevents the use of rigid, pre-programmed trajectories and requires a controller that can handle unmodeled geometric constraints.
In the ‘trivial’ task of cutting in a kitchen, the knife repeatedly comes into contact with objects of varying material properties, ranging from soft fruits, such as strawberries, to stiffer objects, such as pumpkins and a rigid cutting board. The soft mechanics of food make accurate simulation challenging and this, in turn, complicates the application of learning- and optimization-based methods. In practical settings, learning approaches require a large number of interactions to converge to an effective control policy Ian Lenz and Saxena (2015); Mitsioni et al. (2019, 2021); Beltran-Hernandez et al. (2024), which not only produces significant food waste but also poses safety risks when experimenting with a knife.
Figure 1: Robotic arm with knife ridigly attached, and snapshots of robotic cutting experiments.
In addition, for dynamic cutting techniques such as the rocking chop, where the blade rolls across the board, classical trajectory tracking and optimization often fail to yield satisfactory results. Executing such motions via predefined trajectories requires precise kinematic models of the blade’s geometry Mu et al. (2023). Furthermore, it demands an exact characterization of the environmental constraints, specifically the height and planarity of the cutting board, as well as the evolving geometry of the food. In contrast, it would be highly desirable to have a robust approach that does not depend on precise information, such as the knife geometry. Conventional trajectory generation methods lack this generalization capability and typically rely on accurate vision sensing for re-planning Long et al. (2013, 2014); Wright et al. (2024).
Impedance control Hogan (1985a) is the standard paradigm for tasks involving stiff and uncertain contacts. However, identifying a reliable impedance for dynamic cutting remains a significant challenge. The system must be sufficiently stiff to penetrate the food, maintain a rectilinear cutting plane, and ensure accurate rolling kinematics. Conversely, it must remain compliant enough to guarantee reliable contact with the board, and moderate the contact forces and prevent jamming in the presence of geometric uncertainties. Whether these competing requirements can be reconciled with a static, uniform impedance remains an open question. Our experimental findings suggest that impedance must instead be modulated according to the specific geometry and phase of the cutting cycle. However, to the best of our knowledge, a formal theory for phase-dependent or adaptive impedance design in the context of robotic cutting with geometric uncertainties has yet to be established.
In this paper, we propose an approach that reduces dependence on accurate modelling information while enabling the generation of complex motions for contact-rich manipulation. Our framework is rooted in the classical methodology of Virtual Model Control (VMC), originally developed to manage the stiff, complex interactions inherent in robotic locomotion Pratt et al. (2001); Larby and Forni (2025). We introduce a novel virtual model controller, comprising a network of virtual springs, dampers, and linkages, to establish robust rolling kinematics for the knife. Rather than relying on pre-programmed trajectories, the desired behavior emerges from the dynamic interaction between the robot’s physical embodiment, the virtual mechanism, and the environment. Methodologically, our approach diverges from existing VMC literature by incorporating a regulated injection of energy, resulting in a stable, periodic rock chop motion. This is achieved via controlled switching between two virtual mechanical topologies that correspond to the rising and falling phases of the knife.
Our design of the virtual mechanisms for cutting is based on intuitive physical choices, which can be described graphically in a simple way. Our virtual model controller does not require extensive training data, and parameter tuning is efficient. The controller modulates the robot impedance in a complex, nonlinear fashion throughout the cutting cycle, driven by real-time interaction features. By leveraging its virtual components, the virtual mechanisms realize the necessary dynamics without requiring explicit trajectory or force tracking, and demonstrates robustness to environmental variations; the cutting action remains viable even if knife, workspace, or object geometry change, albeit with graceful performance degradation. Finally, the integration of sensory feedback enables online monitoring and optimization of the virtual parameters, facilitating the refinement of cutting frequency and force to ensure successful material separation.
The central finding of this paper is that Virtual Model Control (VMC) provides a robust design framework for complex manipulation, significantly reducing the reliance on high-fidelity modeling and the associated computational overhead. The specific contributions of this work are:
  • A novel VMC architecture for autonomous rocking chop motions without explicit trajectory planning;
  • The introduction of regulated energy injection into the VMC framework to sustain periodic motion via controlled mechanical switching;
  • Analytical and experimental demonstrations verifying the stability and periodicity of the cutting motion;
  • Experimental characterization of cutting performance across various types of vegetables, for different thicknesses and cutting speeds;
  • Experimental validation of the controller’s robustness against perturbations of knife geometry and cutting board height;
  • Demonstration of the controller’s portability (retargeting) through implementation on two distinct robotic platforms: the Franka FR3 and the RT Corp Scirus 17;
  • A preliminary study of the online adaptation capabilities enabled by the proposed VMC design.
The paper is organized as follows. A brief overview of the robotic cutting landscape is summarized in Related works. The methodology of virtual model control and a detailed description of the virtual model controller for cutting are provided in Virtual cutting mechanisms design. An energy-based analysis of the controlled robot behavior can be found in Energy-based analysis. Experimental platform and cutting motion provides the details of the implementation and an illustration of the robot motion. Experimental benchmarks are discussed in Experimental result and discussion, with a focus on performance and robustness. Conclusions and further research directions are explored in Conclusion and future works, which also provides a detailed comparison with related works.
2 Related works
Predicting the mechanical interaction during food cutting is inherently challenging. One of the reasons is that developing an accurate physical model of foodstuffs is complex. The force required to cut through a food item depends on numerous factors, including ripeness, skin presence, orientation, slice thickness, and the knife’s rake angle and sharpness Vincent and Lillford (1991); Khan and Vincent (1991); Atkins (2009b, a). Simply applying a high cutting force can yield imprecise cuts or even damage delicate produce such as tomatoes. In earlier studies, cutting force was often treated as an external disturbance, with adaptive controllers employing position and velocity feedback to maintain predetermined trajectories Zeng and Hemami (1997). Impedance control was introduced to regulate cutting force directly, augmented by adaptive laws to accommodate time-varying desired positions and environments Jung and Hsia (1999).
Robotic cutting is commonly approached as a motion tracking problem, often informed by visual sensing. A key challenge lies in formulating the cutting trajectory itself. While basic trajectories can be designed, more sophisticated cutting trajectories are harder to derive. Remarkably, Long et al. (2013) obtained a smoothed path from camera data to separate meat muscle. Long et al. (2014) proposed a cut and slice trajectory adjusted through force sensing, to reduce damage to deformable objects. More complex rolling (or rocking) actions have been addressed by leveraging the knife geometry to maintain a constant contact force Mu et al. (2023). Straižys et al. (2023) explored learning from demonstration, introducing a coupling term via the Udwadia–Kalaba method to suppress lateral blade movements. Reinforcement learning is used in Beltran-Hernandez et al. (2024) to learn a cutting trajectory and impedance parameters. Because developing an exact physical model of foodstuffs is challenging, learning-based approaches have also garnered attention. A series of works Ian Lenz and Saxena (2015); Mitsioni et al. (2019, 2021) employs data-driven model predictive control (MPC), where a neural network is used to predict future knife positions based on current position and force measurements, thus enabling real-time optimization of the reference force. In Rezaei-Shoshtari et al. (2020), variational autoencoders are used to model the cutting dynamics, though their primary goal is to detect material properties and thickness rather than to learn a cutting strategy.
The main difference with our approach is that our controller can generate complex rocking motion without requiring a predefined trajectory. Our controller design and deployment require minimal information about the food and the external environment (knife geometry, cutting board features). Furthermore, its design is not data-intensive and both design and deployment require minimal computing efforts. In turn, these features guarantee the robot achieves successful cutting, robustly handling different food types.
A second challenge involves managing unanticipated interactions with the environment. Inaccurate camera-based detection of the object position, or variations in the table height, can lead to collisions or excessive forces if the robot strictly follows a fixed cutting path. This often results in safety triggers when the commanded position is physically unachievable (e.g., below the cutting board). To deal with the interaction, Yan et al. Mu et al. (2019, 2023) adopt a hybrid scheme that switches to impedance or direct force control upon detecting contact with the cutting surface. Their method segments the cutting process into distinct press, push, and slice phases, applying different controllers in each stage. In our approach, the controller does not require the identification of different stages of cutting. Rocking is the result of the continuous interaction between robot mechanics, virtual model controller, and environment. Unanticipated interactions may degrade the performance, but only large perturbations will disrupt the whole rocking motion.
A consistent benchmarking framework remains largely absent in the field of robotic cutting. Although prior work often reports end-effector trajectory and force sensor data for individual cutting trials, these metrics appear insufficient for real-world applications. Reliability, that is, the ability to repeatedly produce separated slices of accurate thickness and handle variations in the environment (e.g., knife, board height, or food properties), remains understudied. In contrast, in this paper we quantify the performance of our controller against a broad set of relevant metrics, such as cut frequency, cutting force, cutting speed, slice-thickness accuracy and variance, and we demonstrate successful cuts across different environmental conditions.
3 Virtual cutting mechanisms design
3.1 Virtual model control in a nutshell
A virtual model controller consists of a (virtual) mechanical linkage combined with springs, dampers, and masses that, once coupled with the robot, shapes its behavior. It achieves this by imposing constraints on the motion of the robot and its reaction to the environment. The associated virtual forces are then realized through the robot’s available actuation, typically mapped into joint torques and executed by the motors at each joint Pratt et al. (2001); Zhang et al. (2024); Larby and Forni (2025). This mapping relies on the robot’s kinematics.
For fully actuated robot manipulator based on revolute joints, using $\boldsymbol{q}$ to denote the joint coordinates of the robot, the force $\boldsymbol{f}_{i}$ generated by a virtual component at a point $\boldsymbol{p}_{i}=\boldsymbol{h}_{i}(\boldsymbol{q})$ of the robot is mapped into corresponding joint torques $\boldsymbol{\tau}$ using the Jacobian matrix $J_{i}(\boldsymbol{q})=\frac{\partial\boldsymbol{h}_{i}(\boldsymbol{q})}{\partial\boldsymbol{q}}$. From the principle of virtual work Spong et al. (2005), the torque command to emulate the forces of the virtual model controller is given by
$\boldsymbol{\tau}=J_{1}(\boldsymbol{q})^{T}\boldsymbol{f}_{1}+J_{2}(\boldsymbol{q})^{T}\boldsymbol{f}_{2}+J_{3}(\boldsymbol{q})^{T}\boldsymbol{f}_{3}+\dots$ (1)
Virtual model control is intrinsically stable, because the controlled robot is the interconnection of the passive mechanics of the robot and the passive virtual mechanics Larby and Forni (2025). Stability is guaranteed by passivity theory Willems (1972); van der Schaft (1999); Secchi et al. (2007); Stramigioli (2015); Folkertsma and Stramigioli (2017); Rashad et al. (2020); Ortega et al. (2021). Virtual model control provides an intuitive, task-oriented method for energy shaping and damping injection in robotics Ortega et al. (2001, 1998).
The selection of spring and damper parameters and their geometric configuration also shape the impedance of the robot Hogan (1985a); Colgate and Hogan (1988). Although impedance control is typically implemented with linear springs and dampers, nonlinear springs featuring force saturation can limit excessively rapid movements and enhance safety. In this work, all springs in the virtual mechanism have a saturating nonlinear characteristic between displacement and force given by
$\boldsymbol{f}(k,\sigma,\boldsymbol{z})=\sigma\tanh(k|\boldsymbol{z}|/\sigma)\frac{\boldsymbol{z}}{|\boldsymbol{z}|},$ (2)
where $k$ is the stiffness parameter, the maximum force is governed by $\sigma\geq 0$, and $\boldsymbol{z}$ represents the spring displacement. Fig. 2 shows the force profile of a spring with stiffness 25 N/m and maximum force 20 N.
Figure 2: 2020 N saturated nonlinear spring force profile.
3.2 Virtual model control for cutting
(a) Rocking motion overview. The knife pivots from tip to heel against the cutting board.
(b) Left: Maintaining contact. A spring and damper link the knife tip $\boldsymbol{p}_{1}$ to $\boldsymbol{p}_{a}$, keeping the blade in contact with the board while being able to slide along the rocking direction ($y$ axis). Middle and right: Cutting and raising. Spring-damper pair connects the knife handle $\boldsymbol{p}_{2}$ to $\boldsymbol{p}_{b}$ which moves vertically for cutting and raising the knife. When switch from cutting to raising, the reference points are shifted along $x$ axis by one slice thickness to produce multiple slices.
(c) Switching references smoothly for cutting (left) or raising (right). Spring-damper pair connects $\boldsymbol{p}_{b}$ to $\boldsymbol{r}_{2,1}$ for cutting. Switching from $\boldsymbol{r}_{2,1}$ to $\boldsymbol{r}_{2,2}$ raises the knife. Instead of directly connecting $\boldsymbol{p}_{2}$ to $\boldsymbol{r}_{2,1}$ or $\boldsymbol{r}_{2,2}$, the mass-spring-damper here produces a smoother transition in $z$ axis. $\boldsymbol{p}_{b}$ slides along a rail. The rail is rigidly connected to $\boldsymbol{r}_{2,1}$ and $\boldsymbol{r}_{2,2}$.
(d) Constraining the slicing plane. Left: Define four fixed points $p_{3},\dots,p_{6}$ relative to the knife. Middle: Define four virtual masses, each can slide along the two axes in the desired slicing plane. Right: Each of the points $p_{3},\dots,p_{6}$ is connected to an individual point mass with a spring and damper.
Figure 3: Virtual mechanism design for rocking motion.
A rocking motion involves continuously shifting the knife’s contact point along the cutting board, from its tip to its heel, as illustrated in Fig. 3(a). To achieve this using a virtual mechanism, we first define a virtual mass that slides along the rocking direction ($y$ axis in Fig. 3(b)) and attach a spring and damper from the knife tip at position $\boldsymbol{p}_{1}$ to the position of the virtual mass $\boldsymbol{p}_{a}$ on the cutting board, ensuring that the blade remains in contact with the surface. The goal of the free-sliding virtual mass is to ensure that the knife can move freely while keeping contact with the board, without interfering with the rocking motion. The virtual mass’s position in the slicing direction $y_{\boldsymbol{p}_{a}}$ follows the mass-spring-damper dynamics:
$m_{1}\ddot{y}_{\boldsymbol{p}_{a}}+\boldsymbol{f}(k_{1},\sigma_{1},y_{\boldsymbol{p}_{a}}-y_{\boldsymbol{p}_{1}})+c_{1}(\dot{y}_{\boldsymbol{p}_{a}}-\dot{y}_{\boldsymbol{p}_{1}})=0$ (3)
where $y_{\boldsymbol{p}_{1}}$ is the position of $\boldsymbol{p}_{1}$ in the slicing direction. $c_{1}$ is the damping coefficient.
Another spring and damper connect the knife handle at $\boldsymbol{p}_{2}$ to the control point $\boldsymbol{p}_{b}$ for the cutting motion. Switching $\boldsymbol{p}_{b}$ from reference $\boldsymbol{r}_{2,1}$ to reference $\boldsymbol{r}_{2,2}$ raises the knife. Switching from $\boldsymbol{r}_{2,2}$ to $\boldsymbol{r}_{2,1}$ cuts down (Fig. 3(b)). Hence, the system alternates between a cutting state and a knife-raising state, producing the rocking action. The motion is in cutting phase when $\boldsymbol{p}_{b}$ is connected to $\boldsymbol{r}_{2,1}$ and raising phase when connected to $\boldsymbol{r}_{2,2}$.
Positions $\boldsymbol{p}_{1}$ and $\boldsymbol{p}_{2}$ are defined relative to the robot’s end-effector, assuming that the knife grip is rigid enough, so that the relationship between end-effector and knife does not change. $\boldsymbol{p}_{1}$ and $\boldsymbol{p}_{2}$ do not have to be on the knife. As long as $\boldsymbol{p}_{1}$ is towards the direction of knife tip and $\boldsymbol{p}_{2}$ is in the direction of handle, the motion is feasible. $\boldsymbol{p}_{a}$, $\boldsymbol{r}_{2,1}$ and $\boldsymbol{r}_{2,2}$ lie in the world frame and depend on the position of the food. The difference in $y$ coordinate of $\boldsymbol{r}_{2,1}$ and $\boldsymbol{r}_{2,2}$ regulates the slicing distance. The cutting force is regulated through the stiffness parameters $k_{1}$ and $k_{2}$. The position of $\boldsymbol{p}_{a}$ and the selection of the reference $\boldsymbol{r}_{2,1}$ also affect the cutting force.
The switch from cutting to raising is triggered by the knife’s orientation. When $z$ coordinates of $\boldsymbol{p}_{1}$ and $\boldsymbol{p}_{2}$ are close (within tolerance $\delta_{1}$), the knife spine (the line that passes $\boldsymbol{p}_{1}$ and $\boldsymbol{p}_{2}$) aligns with the cutting board surface and the cut is done. The switch from raising to cutting is triggered when $\boldsymbol{p}_{2}$ is close to $\boldsymbol{r}_{2,2}$ (within tolerance $\delta_{2}$ measured along the z coordinates), to ensure enough height to cut the food. When knife is moving fast, tolerances $\delta_{1}$ and $\delta_{2}$ need to be set large enough to prevent excessive raising or rolling of the knife.
All together, these components produce virtual forces
$\boldsymbol{f}_{1}$ $=\boldsymbol{f}(k_{1},\sigma_{1},\boldsymbol{p}_{a}-\boldsymbol{p}_{1})+c_{1}(\dot{\boldsymbol{p}}_{a}-\dot{\boldsymbol{p}}_{1})$ (4a)
$\boldsymbol{f}_{2}$ $=\boldsymbol{f}(k_{2},\sigma_{2},\boldsymbol{p}_{b}-\boldsymbol{p}_{2})+c_{2}(\dot{\boldsymbol{p}}_{b}-\dot{\boldsymbol{p}}_{2}).$ (4b)
where $\boldsymbol{p}_{b}$ is switched between $\boldsymbol{r}_{2,1}$ and $\boldsymbol{r}_{2,2}$ depending on whether it is in the cutting or raising phase. The way $\boldsymbol{p}_{b}$ switches is also modulated by the dynamics in Fig. 3(c). We smooth the transition by placing a mass that slides along the line between $\boldsymbol{r}_{2,1}$ and $\boldsymbol{r}_{2,2}$. The mass is rigidly attached to $\boldsymbol{p}_{b}$ and connected to $\boldsymbol{r}_{2}$ by spring and dampers.
Finally, to ensure the knife moves primarily within the plane that defines the slice, we introduce four additional points $\boldsymbol{p}_{3},\dots,\boldsymbol{p}_{6}$ on the knife (Fig. 3(d)) and connect each to a virtual mass free to slide along the desired slicing plane, whose position is denoted by the reference $\boldsymbol{r}_{S}$. Each pair is linked by its own spring and damper. This virtual mechanism enforces a plane-constrained motion, yielding a cleaner cut. Attachment points $\boldsymbol{p}_{3}$$\boldsymbol{p}_{6}$ are defined relative to the knife plane with respect to the end effector. These points need not lie on the knife itself; increasing their distance from the blade enhances orientation correction. While three points are theoretically sufficient to constrain orientation, using four provides improved stability without further increasing stiffness. To produce multiple slices in sequence, once a slice is completed and the knife has returned to its raised position, we shift the slicing plane $\boldsymbol{r}_{S}$, $\boldsymbol{p}_{a}$, $\boldsymbol{r}_{2,1}$, and $\boldsymbol{r}_{2,2}$ to the next desired location.
4 Energy-based analysis
Energy considerations provide insight into the emerging periodic motion. The mechanical energy of the robot, $E_{R}$, satisfies the classical passivity inequality Spong (2022); Spong et al. (2005); Ortega et al. (1998)
$\dot{E}_{\mathrm{R}}\leq\dot{\boldsymbol{q}}^{T}\boldsymbol{\tau}+\sum_{i}\dot{\boldsymbol{s}}_{i}^{T}\boldsymbol{d}_{i}$ (5)
where $\boldsymbol{q}$ is the vector of joint coordinates of the robot manipulator, $\boldsymbol{\tau}$ are the related motor torques, and $(\boldsymbol{s}_{i},\boldsymbol{d}_{i})$ are additional co-located positions/forces variables, accounting for generic external actions to the robot due to the interaction of the knife with food and cutting board. The inequality takes into account unavoidable losses due to internal mechanical dissipation.
The controller is also a (virtual) mechanism, whose mechanical (virtual) energy satisfies
$\dot{E}_{\mathrm{VMC}}=-\sum_{i=1}^{6}\dot{\boldsymbol{p}}_{i}^{T}\boldsymbol{f}_{i}-W_{\mathrm{dissipation}}+W_{\mathrm{elastic}}.$ (6)
$W_{\mathrm{dissipation}}\geq 0$ accounts for the power dissipated in the dampers of the virtual mechanism. $W_{\mathrm{elastic}}$ captures the power flow caused by the switching of the references $\boldsymbol{r}_{2,1}$, $\boldsymbol{r}_{2,2}$, and $\boldsymbol{r}_{S}$. The negative sign of the term $\sum_{i=1}^{6}\dot{\boldsymbol{p}}_{i}\boldsymbol{f}_{i}$ in (6) is due to the mechanical interconnection between robot and virtual model controller. Through (1), each force $\boldsymbol{f}_{i}$ acting on the robot exerts an equal and opposite force on the virtual mechanism.
Taking into account the identity $\dot{\boldsymbol{q}}^{T}\boldsymbol{\tau}=\dot{\boldsymbol{q}}^{T}\sum_{i=1}^{6}J_{i}(\boldsymbol{q})^{T}\boldsymbol{f}_{i}=\sum_{i=1}^{6}\dot{\boldsymbol{p}}_{i}^{T}\boldsymbol{f}_{i}$, the total energy of the controlled robot $E=E_{\mathrm{R}}+E_{\mathrm{VMC}}$ satisfies
$\dot{E}\leq-W_{\mathrm{dissipation}}+W_{\mathrm{elastic}}+\sum_{i}\dot{\boldsymbol{s}}_{i}^{T}\boldsymbol{d}_{i}$ (7)
(7) shows that the energy of the controlled robot must continuously decrease if no external forces act on the robot (i.e. $\boldsymbol{d}_{i}=0$) and if we do not switch references $\boldsymbol{r}_{2,1}$, $\boldsymbol{r}_{2,2}$, and $\boldsymbol{r}_{S}$ of the virtual mechanism. Likewise, the energy necessarily decreases also if the external forces $\boldsymbol{d}_{i}$ represent interaction forces with the food and cutting board. This follows from passivity theory Willems (1972); Ortega et al. (1998); van der Schaft (1999); Secchi et al. (2007), as the complete mechanical system consisting of the controlled robot interacting with food and cutting board can be modelled as the negative feedback interconnection of passive systems. In both scenarios, without switching, the robot motion eventually settles in a resting position. For the gravity-compensated robot, this corresponds to a minimum of the elastic potential energy. This is a classical result of passivity-based control Ortega et al. (2001).
In our case, each time a reference is switched at a specific robot configuration, a bounded amount of energy is injected into or removed from the system. This action continuously drives the controlled robot into motion. The switch between the cutting and raising phases is governed by specific triggers: the transition from raising the knife to cutting is based on the position of the knife, while the switch from cutting to raising is mainly triggered by the orientation of the knife. When a switch occurs, the spring connecting $\boldsymbol{p_{b}}$ to $\boldsymbol{r}_{2,i}$, $i\in\{1,2\}$, is instantly detached from $\boldsymbol{r}_{2,1}$ and attached to $\boldsymbol{r}_{2,2}$ and vice-versa. The injected/removed energy corresponds to the discontinuity in elastic potential energy resulting from the instantaneous change in spring extension.
The periodic cutting motions of the controlled robot, illustrated in the experimental results section below, emerge at the equilibrium between the energy injected via reference switching and the total dissipation from virtual damping, robot internal dissipation, and environmental resistance (i.e. food and cutting board). Along any periodic motions of period $\bar{T}$ the controlled robot must satisfy
$\int_{t}^{t+\bar{T}}\!\!\!\!\!\!\!W_{\mathrm{elastic}}(\tau)d\tau=\!\int_{t}^{t+\bar{T}}\!\!\!\!\!\!\!W_{\mathrm{dissipation}}(\tau)\!+\!\sum_{i}\dot{\boldsymbol{s}}_{i}^{T}(\tau)\boldsymbol{d}_{i}(\tau)d\tau$ (8)
for all times $t$ (where we have neglected the internal dissipation of the robot, for simplicity).
The energy injected at each switch is dissipated by the robot’s inherent damping and the virtual dampers, or transferred to the environment through interaction with the food and the cutting board. Any energy surplus ultimately manifests as increased velocity or higher cutting forces, both of which enhance dissipation. Due to the geometry of the compliant constraints imposed on the knife by the virtual mechanisms, the switch-induced variations primarily affect the knife’s planar velocity and its contact forces. Hence, over multiple cycles, the system settles into a stable periodic regime.
Further insights can be obtained by examining a simplified single-degree-of-freedom model, representing a first-order approximation of the vertical dynamics of $\boldsymbol{p}_{2}$
$m\ddot{z}+c\dot{z}+k(z-r)=0,$ (9)
where $m>0$ is the mass, $c>0$ is the damping coefficient, and $k>0$ is the spring stiffness. $r$ is the reference, switching between two values, $r\in\{r_{1},r_{2}\}$. The mechanical energy reads
$E=\frac{1}{2}m\dot{z}^{2}+\frac{1}{2}k(z-r)^{2}$ (10)
from which
$\dot{E}=-c\dot{z}^{2}+W_{r}.$ (11)
$W_{r}$ accounts for the elastic energy injected or removed at switching times (an impulse). This can be computed by looking at the left and right limits of the switching time $t$, denoted by $t^{-}$ and $t^{+}$,
${E(t^{+})\!-\!E(t^{-})}=\frac{k{(z(t^{+}\!)-r(t^{+}\!))}^{2}}{2}-\frac{k{(z(t^{-}\!)-r(t^{-}\!))}^{2}}{2}.$ (12)
On a period, injection of energy and dissipation must balance. An example is provided in Fig. 4, based on a switching law that captures the fundamental transitions between cutting and raising in the rocking motion:
$r$ $=r_{2}\quad\text{if}\quad\bigl(\dot{z}\geq 0\,\text{and}\,z\leq z_{2}\bigr)\;\text{or}\;\bigl(\dot{z}\leq 0\,\text{and}\,z\leq z_{1}\bigr),$
$r$ $=r_{1}\quad\text{if}\quad\bigl(\dot{z}\geq 0\,\text{and}\,z\geq z_{2}\bigr)\;\text{or}\;\bigl(\dot{z}\leq 0\,\text{and}\,z\geq z_{1}\bigr).$
The phase plane of Fig. 4 shows that solutions converge to a closed curve. Intuitively, if the mass moves too fast, the damper dissipates more energy, returning the system toward this closed curve. Conversely, if motion is too slow, less energy is dissipated, and the system gains more momentum until it approaches the same closed curve.
Figure 4: Phase plane of a mass spring damper system with switching references $r_{1}=-1$ and $r_{2}=1$ and positions $z_{1}=-0.8$ and $z_{2}=0.8$. Trajectories with different initial conditions converge to the same limit cycle.
To conclude this section, we provide a practical estimation of the robot’s work per cycle. At steady state, this metric quantifies the energetic cost of the cutting process itself (interaction with food and the environment) by isolating it from internal mechanical and virtual damping losses.
$E_{\mathcal{C}}=\int_{t}^{t+\bar{T}}\boldsymbol{F}(\tau)^{T}\dot{\boldsymbol{x}}(\tau)d\tau,$ (13)
where $\bar{T}$ is the cycle period, $\boldsymbol{F}$ is the force at the robot end-effector, and $\boldsymbol{x}$ its position. Practically this force is sampled by a force sensor, leading to the approximated formula
$E_{\mathcal{C}_{j}}\ =\!\!\!\!\sum_{t_{i}\in\mathcal{C}_{j}\setminus\{t_{N}\}}\boldsymbol{F}^{T}\bigl(t_{i}\bigr)\bigl(\boldsymbol{x}(t_{i+1})-\boldsymbol{x}(t_{i})\bigr),$ (14)
where $\mathcal{C}_{j}=\{t_{1},\dots,t_{N}\}$ contains the sampling time instants of cycle $j$ at which sample measures are taken.
Fig. 5 illustrates the work done during a carrot-cutting experiment. Initially, the knife performs a pure rocking motion without food. Between the 7th and 16th cycles, a carrot is introduced and processed, after which it is removed. Fig. 5 provide indirect evidence of the stability of the achieved periodic motion. The work done increases due to the additional dissipation of cutting through the carrot but the system returns to its baseline energy level once the carrot is removed. This behavior demonstrates the virtual model controller’s ability to compensate for external perturbations and consistently restore the original energy balance.
Figure 5: Work done for each cutting cycle. Red: cutting a carrot, blue: no food
5 Experimental platform and cutting motion
5.1 Robot platform and control architecture
The virtual model controller outlined in the previous section is implemented through the VMRobotControl.jl package Larby (2024), a platform-independent framework designed for virtual model control in robotic systems. This package enables the construction and simulation of both the robot and virtual components. Given the robot’s URDF representation, the package offers several primitives to define the virtual model controller, to compute its dynamics, and to generate the torque commands for each joint.
Our main experimental platform consists of a 7-DoF Franka Research 3 arm, with a maximum payload of 3 kg. To collect force data, a ROBOTIQ FT 300-S force-torque sensor is mounted at the robot wrist, providing force and torque measurements at 100 Hz with a noise level of approximately 0.1 N. A knife is rigidly connected to the sensor using a 3D-printed attachment as shown in Fig. 1. The portability of the virtual model controller allowed us to implement and test it also on the humanoid robot Sciurus17, a lightweight and low-cost platform developed by RT Corp. Scirus17 features 17 degrees of freedom (DoF) distributed across its upper body (with 7 DoF per arm, 2 DoF for the head and 1 DoF for the torso). The robot utilizes Dynamixel motors with current control, and the conversion from torque to current is approximated by means of a linear relation based on the performance characteristics of the motor.
The controller runs on a standard computer. The overall control architecture is depicted in Fig. 6. Virtual forces $\boldsymbol{f}$ generated by the virtual mechanisms are computed and translated into joint torques $\boldsymbol{\tau}$ using (1). The computation of the current state of the virtual mechanisms, such as the elongation of virtual springs and the velocities used by the virtual dampers, is based solely on the robot’s joint position $\boldsymbol{q}$ and velocity $\dot{\boldsymbol{q}}$. All force computations and their conversion into joint torques rely on forward kinematics and the associated Jacobians. This represents a significant departure from traditional motion planning, where execution is typically based on inverse kinematics.
Figure 6: Control architecture implemented with VMRobotControl.jl package.
5.2 The cutting motion in experiments
(a) Snapshots from experiment: rocking a courgette with a knife.
(b) Top: Solid: $\boldsymbol{p}_{1}$, Dashed: $\boldsymbol{p}_{a}$. Bottom: Solid: $\boldsymbol{p}_{2}$, Dashed: $\boldsymbol{p}_{b}$.
(c) End effector position and force reading.
(d) End effector orientation in Euler angles and torque reading.
Figure 7: Rocking motion as a result of the interaction between the virtual mechanism and environment.
The virtual mechanisms constrain the motion of the blade by enforcing compliance at contact points and ‘soft’ constraints. Switching of the reference at specific points induces a cyclical motion of the knife, which is determined by the interaction among robot dynamics, compliant virtual mechanism constraints, and contacts with the environment. A first illustration is provided in Fig. 7, whose control parameters are specified in Table 1 (a). Using courgette as the cutting target, the snapshots in Fig. 7(a) illustrate the continuous contact between the knife and the cutting board while the knife rolls back and forth. The successful separation of the courgette is also seen in the snapshots. Since the motion is not pre-programmed, variations in the knife geometry or cutting board features will yield different motions.
Fig. 7(b) shows the position of two springs responsible for the rocking motion. The sliding of the virtual mass ensures that $\boldsymbol{p}_{a}$ stays close to $\boldsymbol{p}_{1}$ and no additional force is applied in the rocking direction. The displacement between $\boldsymbol{p}_{1}$ and $\boldsymbol{p}_{a}$ along the $z$ axis pulls the knife towards the cutting board and ensures contact. The knife moves to the next slice position when raised. Fig. 7(c) and 7(d) show the end-effector’s position and force, orientation and torque. The resulting end-effector’s motion is a fairly complicated trajectory, difficult to design manually. One would need to know the geometry of the knife to decide where the end-effector should be while rolling the knife back and forth.
Force and torque are measured at the robot’s wrist and transformed to the world coordinate. $F_{y}$ and $F_{z}$ correspond to the forces applied to the robot wrist to slice and cut. $F_{x}$, $\tau_{y}$ and $\tau_{z}$ act to stabilize the knife on the slicing plane. The pitch and yaw remain small, which shows that the stabilization of the slicing plane is effective. The roll and torque $\tau_{x}$ account for the rocking motion. During the lifting of the knife, a higher amount of torque $\tau_{x}$ is applied when the contact point is shifted further to the tip.
6 Experimental result and discussion
6.1 Forces and velocities
(a) Black: Force reading in the z axis. Red: Switch from raising to cutting. Blue: Switch from cutting to raising. Dash: Same control parameters for all food. Continuous: Tuned control parameters for each food so that the knife cuts through without applying excessive force.
Figure 8: Rocking mechanism applied to five different food. Left two columns: blue - untuned, red - tuned for minimal force.
To evaluate the controller’s versatility, we conducted cutting experiments on a variety of vegetables (spring onion, cucumber, carrot, courgette, and potato) under two conditions: (i) a uniform parameter set applied to all food types, and (ii) food-tuned parameters manually optimized to minimize contact force. Reducing the cutting force helps to achieve precise cuts and prolongs the lifespan of both the knife and the cutting board. Larger items, such as cucumber, courgette, and potato, were cut in half to make them suitable for rocking and ensure stable placement on the cutting board. All vegetables were securely held to prevent motion during cutting. Results are summarized in Fig. 8. The rocking motion was initially validated through simulation but parameter tuning was performed on the physical robot to account for the complexities of modeling contact forces during cutting. Although we explore the influence of these parameters on performance, a systematic optimization study remains beyond the scope of this work. We refer the reader to Larby and Forni (2025) for an optimization framework tailored to virtual model control.
To quantitatively evaluate the cutting process, we have used the following metrics. Let $\mathcal{C}_{\mathrm{cut},i}=\{t_{1},\dots,t_{N_{i}}\}$ denote the set of sampling times during the cutting phase $i$. These belong to the time intervals in which $\boldsymbol{p}_{b}$ is connected to the reference $\boldsymbol{r}_{2,1}$, as shown in Fig. 3(c) (left). The average and peak forces during the $i$th cycle are computed from the force-torque sensor reading, mounted at the robot wrist
$F_{\text{avg},i}\,=\!\!\sum_{t\,\in\,\mathcal{C}_{\mathrm{cut},i}}\!\!\frac{|\boldsymbol{F}(t)|}{N_{i}}\qquad\quad F_{\text{peak},i}\,=\max_{t\,\in\,\mathcal{C}_{\mathrm{cut},i}}|\boldsymbol{F}(t)|.$ (15)
Similarly, the average and maximum velocities of the robot wrist during cutting are estimated as follows.
$v_{\text{avg},i}\,=\!\!\sum_{t\,\in\,\mathcal{C}_{\mathrm{cut},i}}\!\!\frac{|\boldsymbol{v}(t)|}{N_{i}}\qquad\quad v_{\text{peak},i}\,=\max_{t\,\in\,\mathcal{C}_{\mathrm{cut},i}}|\boldsymbol{v}(t)|,$ (16)
where $\boldsymbol{v}$ is the velocity of the wrist, computed using the robot kinematics. Finally, cut frequency is defined as
$f_{\text{cut}}\,=\,\frac{N_{\text{switch}}-1}{T},$ (17)
where $N_{\text{switch}}$ is the total number of transitions from the cut to the raise state that occur during the time interval $T$ of a whole experiment. This frequency indicates how rapidly consecutive slices are produced. Work for one slice is computed with (14).
When using a single set of control parameters for all foods, the force profiles in the $z$-axis (Fig. 8(a)) were similar across all food types despite significant differences in the material properties of different food. This is because the control parameters were chosen to apply a sufficiently high force so that the intrinsic properties of the food had minimal impact on the resulting behavior. Peak force is typically observed near the transition from cut to raise and varies with the energy absorption characteristics of each vegetable. Higher levels of ‘residual’ energy after cut typically results in a higher contact forces before the knife is retracted. With parameters tuned specifically for each food type, the force required for cutting spring onion and cucumber is noticeably reduced, while high force levels are needed for other vegetables. In terms of tuning strategy, stiffness of springs was uniformly reduced (increased) to reduce (increase) contact forces. The mass–spring–damper parameters ($m_{b},k_{b},c_{b}$) were adjusted accordingly to preserve cutting frequency. For smaller items, the knife’s lifting height was reduced by increasing the switching tolerance $\delta_{2}$, maintaining comparable cutting frequencies even with lower cutting velocities. For larger items, the reference position for raising the knife, $\boldsymbol{r}_{2,2}$, was set higher to accelerate the upward motion, with $\delta_{2}$ adjusted accordingly. Compared to the uniform parameter, cut frequency remains similar or becomes higher while velocity during cutting is reduced for carrot, courgette and potato.
These results do not provide an exhaustive characterization, which is beyond the scope of the current paper, as this would require a proper optimization framework and extensive validation tests for a proper statistical analysis, further complicated by the extreme variability in mechanical properties of vegetables, Agrawal and Lucas (2003). However, they illustrate the tunability of the virtual model controller, by showing how different parameter configurations can achieve successful cutting while producing distinct motion characteristics, allowing the control designer to shape the desired behavior through physical parameter selection rather than trajectory specification. Fig. 1 shows the sliced vegetables at the end of experiments. By visual inspection, both (i) uniform and (ii) food-specific parameter selection produce no noticeable difference in slice quality or accuracy.
Figure 9: Cutting different slices. Top row: blue - 0.94 cut/sec, red - 0.51 cut/sec. Bottom row: fast and slow frequencies.
(a) Rocking with the proposed VMC succeeds for knives with different sizes and blade profiles. In contrast, impedance control following a pre-defined trajectory fails under variations in knife geometry. The reference trajectory is the end-effector position and orientation recorded during a successful cut with Knife 1 under VMC. Trajectories of the knife tip, heel and the end-effector are shown in each figure.
(b) Forces, velocities, and cut frequencies for different knives. Blue: cutting with the proposed VMC; red: impedance control on predefined trajectory.
Figure 10: Rocking motion with different knives using the same virtual mechanism and control parameters.
6.2 Accuracy
To evaluate the precision of our cutting mechanism, we cut courgette samples using various target thicknesses and cutting speeds. Two average cut frequencies were implemented for each thickness setting, approximately 0.51 and 0.94 cuts per second. Fig. 9 shows representative slices obtained under these conditions.
Cutting accuracy is assessed by measuring each cut in four distinct locations. Measurements are taken with care, avoiding deformation of the slice as much as possible. We use three metrics to measure (i) if the desired thickness is achieved, (ii) if the accuracy is consistently achieved with continuous cutting, and (iii) if each slice is evenly cut.
For (i), we compare the average of all measurements $\Delta$ with the desired one. We compute
$\Delta=\frac{1}{N}\sum_{i=1}^{N}\Delta_{i}\qquad\quad\Delta_{i}=\frac{1}{4}\sum_{j=1}^{4}\Delta_{i,j}$ (18)
where $N$ is the total number of slices, $\Delta_{i}$ is the average thickness of slice $i$ based on four thickness measurements $\Delta_{i,j}$. For (ii), we take the variance of the average thickness slices. Namely,
$\sigma^{2}_{\Delta}=\frac{1}{N}\sum_{i=1}^{N}\left(\Delta_{i}-\Delta\right)^{2}.$ (19)
Higher variance indicates lower consistency. Finally, for (iii), we take the variance of the four thickness measurements within each slice and compute its average across all slices:
$\sigma^{2}_{\Delta}=\frac{1}{N}\sum_{i=1}^{N}\sigma^{2}_{\Delta_{i}}\qquad\sigma^{2}_{\Delta_{i}}=\frac{1}{4}\sum_{j=1}^{4}\left(\Delta_{i,j}-\Delta_{i}\right)^{2}.$ (20)
In this case, a higher variance indicates lower evenness.
Fig. 9 shows that average thickness is close to the desired target. In particular, the slices are consistently slightly thinner than expected. Several factors may contribute to this discrepancy: the knife’s movement matches the desired thickness without compensating for its own width, alongside the plastic deformation of the vegetable during cutting and moisture loss in the slices. Despite these effects, achieving nearly one slice per second with such precision is comparable to or even exceeds human performance.
Faster cutting generally degrades performance in terms of average thickness, consistency, and evenness, as illustrated in Fig. 9, except in the case of 1-mm slices. The decreased performance at higher speeds is mainly due to the limited ability of the knife to accurately reach the target slicing position in a short time frame, resulting in fluctuations and positional adjustments during the cut. This effect is minimized when cutting very thin slices. Additionally, 1-mm slices experience greater plastic deformation at slower cut speeds, potentially hindering the initiation and propagation of a clean cut.
(a) Experiment setup with different cutting board heights.
(b) Forces and velocities for different heights of the board. Impedance control fails to complete the motion at a +24 mm height with high gains and at +32 mm with low gains due to excessive contact forces on the end-effector, while the proposed VMC continues to operate. Blue: cutting with the proposed VMC; red: impedance control (high gain); yellow: impedance control (low gain).
Figure 11: Rocking with different cutting boards using the same virtual mechanism and control parameters.
6.3 Robustness
Our virtual model controller demonstrates inherent robustness. As shown in previous section, the same set of control parameters achieved successful cut of a variety of food items, from delicate spring onions to rigid potatoes. In addition to these results, we tested the controller using two different knives. Knife 1 features an optimal blade curvature for rocking and has been used in all experiments so far. Knife 2 has a different blade geometry, which is still suitable for rocking. Results are shown in Fig. 10. To assess robustness, we consider a uniform set of parameters for both knives. Although both knives exhibited rocking motion, the differences in blade geometry affected energy dissipation and resulted in different rocking features. Knife 2 has a smaller blade than Knife 1, resulting in a reduced energy dissipation through contact with the cutting board. This led to a higher residual energy after cutting, which turned into speed, that is, faster cuts. The increase in force peaks for Knife 2 can be explained in a similar way: Knife 2 moves faster than Knife 1, leading to stronger forces around the transition from the cutting phase to the raising phase.
We compared our approach against a standard impedance controller based on pre-defined nominal trajectory. Deriving a suitable rocking trajectory is itself non-trivial, as it requires precise knowledge of the knife geometry and cutting board height. We therefore constructed the nominal trajectory from the end-effector trajectory of a successful cutting motion with Knife 1 generated by the virtual model controller. An impedance controller with high gains (Table 1 (d)) was used for tracking it, tuned manually to achieve the best accuracy without inducing high-frequency oscillations. For impedance control, we define the cutting phase as the time interval during which the end-effector $z$-position is decreasing. Cut forces are then computed using (15). With Knife 1, the rocking motion could be reproduced and cutting was generally successful, although noticeable tracking discrepancies remained. Notably, a thin film of onion at the base was left unseparated in most of the cuts, indicating insufficient normal cutting force. With Knife 2, the geometric mismatch of the knife shape caused the blade to lose contact with the board, leading to poor cuts without complete separation.
We also demonstrate the ability of the virtual model controller to handle different heights of the cutting board. Fig. 11(a) illustrates the experimental setup, where the nominal cutting board height was modified by adding MDF layers underneath. Fig. 11 shows that, without changing any control parameters, the virtual model controller maintained a stable rocking motion across all heights. As the board height increased, the peak cutting force and peak and average cut velocities gradually decreased, which can be explained by the reduced distance available for the knife to accelerate before switching and by the saturating profiles of the nonlinear springs.
Also in this scenario, we compared our approach against an impedance controller, testing high-gain and a low-gain impedance setting (Table 1 (d)). The impedance controller tracked the pre-defined nominal trajectory constructed as specified above. For the cutting board at the nominal height, the low-gain impedance controller failed to cut onions completely, with the cut approximately reaching halfway through due to insufficient applied force. When the cutting board was raised by 8 mm, the same impedance controller was then able to cut through the spring onions, as the higher board effectively increased penetration for a given trajectory. However, further increases in board height led to larger contact forces, triggering protective stops due to excessive forces on the arm. Specifically, the high-gain controller failed at 24 mm height increment, while the low-gain controller tolerated higher displacements but ultimately failed at 32 mm increment.
6.4 Portability
The proposed virtual mechanism is platform independent. We demonstrate this by implementing the same controller on a humanoid robot from RT Corp, Sciurus17. Unlike the Franka platform, Sciurus17 has a lower payload capacity, slower communication speed and uses current control as a proxy of torque control. These differences can be managed by re-tuning the parameters of the controller to ensure smaller forces and slower motions. The cutting plane was kept fixed, and food items were manually fed to the knife to reduce system burden. Fig. 12 shows that the inherent limitations of the robotic platform led to reduced performance. However, successful rocking motion is achieved.
(a) Snapshots of Sciurus17 rocking with a knife.
(b) Top: solid - $\boldsymbol{p}_{1}$, dashed - $\boldsymbol{p}_{a}$. Bottom: solid - $\boldsymbol{p}_{2}$, dashed - $\boldsymbol{p}_{b}$.
Figure 12: Rocking motion Sciurus17. The virtual model controller is unchanged but its parameters have been adjusted.
7 Conclusion and future works
We have proposed a new virtual model controller for rock cutting. The rocking motion has been discussed theoretically, through energy considerations, and experimentally. Performance has been illustrated by cutting several vegetables, for several requirements of thicknesses and cut frequencies. Robustness has been illustrated by using two different knives and several cutting board heights. Figs. 10 and 11 clarify the difference between the virtual model controller and other control methods based on tracking of a nominal, preplanned trajectory. The latter typically require precise knowledge of knife geometry and high-gains for accurate tracking, which limit robustness. In contrast, the rocking motion established by the virtual model controller is not pre-programmed but shaped by the interaction between robot, controller and environment, therefore robustly adapts to a wide range of settings. The experimental evaluation of the impedance controller show the limits of shaping the end-effector impedance without explicitly taking into account the geometry of the task. In contrast, the design of the virtual model controller starts from the task geometry to structure the compliance of the virtual mechanisms in a task-relevant way. This leads to more flexible tuning and increased robustness. For instance, the maximum force $\sigma_{1}$ affects directly the contact force with the cutting board. Likewise, the maximum force $\sigma_{\mathrm{ori}}$ penalizes deviations from the slicing plane. Experiment videos are available in Extension 1.
Successful operation depends on reasonable selection of parameters for the virtual model controller. The position of the object and the orientation of the knife relative to the end effector must be known to ensure effective cutting. However, precise geometric details of the knife or the exact height of the cutting board are not required. Similarly, the exact attachment points of the virtual components are not critical, as approximate placement is sufficient. These features highlight critical differences with Mu et al. (2023), which is the only prior work explicitly considering rolling motion, to the best of our knowledge. In Mu et al. (2023), the motion is decomposed into three stages (pressing, touching, and slicing) combined to different control modes (position, hybrid position–impedance and hybrid position–force control). The implementation requires switching among multiple controllers to track motion and handle contact at different stages of the cut, and fundamentally relies on the explicit geometric model of the knife to plan contact-point transitions and compute associated Jacobians. Mu et al. (2023) demonstrates single-cut performance on selected food items but does not consider repetitive cutting cycles, so robustness under repeated operation and environment variations is not evaluated. In contrast, our virtual-model controller manages both motion generation and interaction without mode switching and operates with substantially less information.
The proposed design is a key component in the development of a fully integrated cutting pipeline. This involves autonomous pick and place of the food, its orientation on the cutting board, and the ability to hold the food in place during cutting. The robot should also be able to recognize different foods and knives, and to measure in real-time the result of its action, during and after cutting. At the level of food-knife interaction, a key future direction is the systematic exploration of adaptation/learning of the control parameters, for stronger performance. The geometric configuration of the virtual model controller serves as the primary driver of the robot’s behavior, while parameters such as stiffness and damping are used for fine-tuning. This provides a fundamental advantage for tuning, as the geometry defines the core motion while the control parameters refine it. The objective is to exploit this structural advantage in combination with external sensing, such as vision and tactile feedback, to enable real-time parameter adaptation. For example, the stiffness $k_{2}$ can be adjusted in real-time to achieve successful food separation while moderating the force exerted by the knife on food and board. The short experimental study proposed below show encouraging preliminary results. Every six cutting cycles we increase $k_{2}$ if the knife does not achieve separation of food and we reduce $k_{2}$ if the knife’s handle overshoot is too large (i.e. if the knife handle moves past the critical reference). This leads to the adaptation law
$k_{2}^{(n+1)}=k_{2}^{(n)}+\alpha\,e,$
where $e$ is the error signal capturing the mismatch between desired and achieved vertical position of the knife, and $\alpha$ is the adaptation gain. Results are summarized in Fig. 13. $k_{2}$ starts from a low value, which is sufficient to cut through a carrot. From time $t=20$s, the carrot presents a harder consistency which requires adaptation. Successful cutting is restored after time $t=35s$ through adaptation of $k_{2}$.
(a) Carrot cut result.
(b) Stiffness $k_{2}$.
Figure 13: Adaptation of spring stiffness for cutting a carrot.
Table 1: Control parameters for each experiment. Units: stiffness $k$ [N/m], maximum force $\sigma$ [N], damping $c$ [N·s/m], position $\boldsymbol{r}$ [m], mass $m$ [kg], $\delta$ [m], $\tau$ [s]. The mass–spring–damper system ($m_{b},k_{b},c_{b}$) generates smooth transitions between reference points. Parameters are selected to ensure overdamped behavior, and the time constant $\tau$ of the slower pole is reported as an intuitive measure of the transition dynamics. Consistent parameters across all experiments: masses for simulating the sliding motion ($m_{a}=m_{\text{ori}}=0.1$). Consistent parameters across all experiments on Franka: $\sigma_{1}=10,c_{1}=1,\sigma_{\text{ori}}=50,c_{\text{ori}}=10$, and $\boldsymbol{r}_{2,1}\,(y,z)=0.43,-0.1$.
(a) Uniform control parameters and parameters tuned to reduce force for individual vegetables. The most sensitive parameters here are the stiffness values ($k_{1}$, $k_{2}$) governing contact and cutting. $m_{b},k_{b},c_{b}$ regulate the transition between reference points $\boldsymbol{r}_{2,1}$ and $\boldsymbol{r}_{2,2}$ to maintain similar cutting frequency when $k_{2}$ is reduced. $\boldsymbol{r}_{2,2}$ and $\delta_{2}$ are adjusted according to object height. Robot: Franka.
Exp $k_{1}$ $k_{2}$ $\sigma_{2}$ $c_{2}$ $k_{\text{ori}}$ $\boldsymbol{r}_{2,2}\,(y,z)$ $\delta_{1}$ $\delta_{2}$ $\tau$
Uniform parameters 25 150 25 2 1200 0.58, 0.4 0.02 0.15 0.5
Spring onion 10 65 25 2 1200 0.58, 0.4 0.02 0.2 0.05
Cucumber 10 75 25 2 1200 0.83, 0.5 0.02 0.25 0.05
Courgette 10 105 25 2 1200 0.83, 0.5 0.02 0.25 0.05
Carrot 20 105 25 2 1200 0.83, 0.5 0.02 0.25 0.05
Potato 20 105 25 2 1200 0.83, 0.5 0.02 0.25 0.05
(b) Different cuttiing frequencies. The primary tuning parameter is $k_{2}$, increased for faster cuts. $m_{b}$, $k_{b}$, and $c_{b}$ are adjusted to switch faster, while $k_{1}$ is increased to ensure stable contact during rapid motion. $\delta_{1}$ is increased for faster cuts to accommodate higher knife velocity and avoid overshoot during motion reversal. Robot: Franka.
Exp $k_{1}$ $k_{2}$ $\sigma_{2}$ $c_{2}$ $k_{\text{ori}}$ $\boldsymbol{r}_{2,2}\,(y,z)$ $\delta_{1}$ $\delta_{2}$ $\tau$
Fast cut (0.94 cut/s) 45 200 30 2 800 0.73, 0.35 0.035 0.145 0.015
Slow cut (0.51 cut/s) 25 120 25 2 800 0.58, 0.4 0.02 0.15 0.5
(c) Robustness to environment variations. Identical mechanism and control parameters are used across all robustness experiments. Robot: Franka.
Exp $k_{1}$ $k_{2}$ $\sigma_{2}$ $c_{2}$ $k_{\text{ori}}$ $\boldsymbol{r}_{2,2}\,(y,z)$ $\delta_{1}$ $\delta_{2}$ $\tau$
Robustness 30 130 25 2 800 0.63, 0.4 0.01 0.2 0.1
(d) Robustness to environment variations. Impedance gains used for the trajectory-tracking baseline. Translational stiffness is given along the Cartesian axes, and rotational stiffness about the corresponding axes. Robot: Franka.
Translational stiffness Rotational stiffness
$K_{x}$ $K_{y}$ $K_{z}$ $K_{r_{x}}$ $K_{r_{y}}$ $K_{r_{z}}$
High gain 1500 7000 7000 300 170 170
Low gain 750 3500 3500 150 85 85
(e) Platform independence. Parameters are adjusted to account for differences in actuation capability and communication frequency between platforms. Stiffness and max forces are reduced. Robot: Sciurus.
$k_{1}$ $\sigma_{1}$ $c_{1}$ $k_{2}$ $\sigma_{2}$ $c_{2}$ $k_{\text{ori}}$ $\sigma_{\text{ori}}$ $c_{\text{ori}}$ $\boldsymbol{r}_{2,1}\,(y,z)$ $\boldsymbol{r}_{2,2}\,(y,z)$ $\delta_{1}$ $\delta_{2}$ $\tau$
10 2 0.3 20 10 0.1 30 10 1 0.4, -0.1 0.6, 0.6 0.047 0.38 0.45
{acks}
Yi Zhang is supported by the Engineering and Physical Sciences Research Council, AgriFoRwArdS Center for Doctoral Training [EP/S023917/1] and RT Corp.